### v1

# Syntomic cohomology of truncated Brown–Peterson spectra

Abstract.
We compute the $\operatorname{MU}$-based syntomic cohomologies, mod $(p,v_{1},\cdots,v_{n})$, of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of the truncated Brown–Peterson spectrum $\operatorname{BP}\langle n\rangle$. As qualitative consequences, we resolve the Lichtenbaum–Quillen, telescope, and redshift questions for the algebraic K-theories of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$. This extends work of the Hahn and Wilson. We also explicitly compute the algebraic K-theory of arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle 2\rangle$ at all primes $p\geq 5$ extending previous work of the author, Ausoni, Culver, Höning, and Rognes.
1. Introduction
By celebrated work of Quillen [QUI69], complex cobordism provides a bridge between geometric topology and algebraic geometry. Working $p$-locally, a homology theory called Brown–Peterson homology $\mathrm{BP}$ splits off of complex cobordism. This homology theory has coefficients a polynomial algebra over the $p$-local integers with generators $v_{i}$ for $i\geq 1$. Complex oriented homology theories carry a formal group law, and the complex oriented Brown–Peterson homology theory carries the universal $p$-typical formal group law. By Lazard [LAZ55], $p$-typical (1-dimensional) formal group laws are classified by their height and we may ask whether there is a complex oriented homology theory that carries a formal group law of height $\leq n$. One way to build such a ring spectrum is to take a quasi-regular quotient of Brown–Peterson homology to produce the truncated Brown-Peterson homology $\mathrm{BP}\langle n\rangle$ with coefficients a polynomial algebra over the $p$-local integers on generators $v_{1}$, $v_{2}$, $\dots$ $v_{n}$.
Structured ring spectra play a fundamental role in homotopy theory, for example more highly structured ring spectra are equipped with a richer notion of power operations. It is known that complex cobordism is a $\operatorname{\mathbb{E}}_{\infty}$-ring, which means it is a homology theory equipped with a coherent commutative multiplication. It was hoped that the Brown–Peterson homology theory and the truncated Brown–Peterson cohomology theories would have similar structure. It was shown that $\operatorname{\mathbb{E}}_{\infty}$ algebra structures exists on $\operatorname{BP}\langle n\rangle$ for n=1,0,1n=-1,0,1 at all primes and $n=2$ at the primes $p=2,3$ by [BR08, HL10, LN12, LN14] and an $\operatorname{\mathbb{E}}_{4}$-algebra structure exists on $\operatorname{BP}$ by [BM13]. However, it was proven by [LAW18, SEN24] that $\operatorname{BP}$ and $\operatorname{BP}\langle n\rangle$ do not have $\operatorname{\mathbb{E}}_{\infty}$ algebra structures when $n\geq 4$, so we do not live in the best of all possible worlds.
Producing $\operatorname{\mathbb{E}}_{m}$ algebra structures on truncated Brown–Peterson spectra for $m>1$ at arbitrary heights $n$ remained an open problem until [HW22] proved that there exist $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra for all primes $p$ and heights $n$. These forms depend on a choice of generators $v_{i}$. The most common choice would be to choose the Araki or Hazewinkle generators. However, it was observed by Strickland [STR99, Remark 6.5] at the prime $2$ that, although these forms of truncated Brown–Peterson spectra are equipped with $\operatorname{\mathbb{E}}_{1}$ algebra structures, they are not equipped with additional structure. In particular, they are not equipped with $\operatorname{\mathbb{E}}_{2}$ algebra structures.
To compute invariants of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra we are therefore forced to work without additional structure. The algebraic K-theory and topological Hochschild homology of an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra is only an $\operatorname{\mathbb{E}}_{0}$ algebra over the corresponding invariant for $\operatorname{MU}$, where an $\operatorname{\mathbb{E}}_{0}$ algebra simply consists of a unit map. Nevertheless, the fact that the unit map $\operatorname{MU}_{(p)}\to\operatorname{BP}\langle n\rangle$ is a $\pi_{*}$-surjection allows us to proceed. We therefore resolve the redshift, Lichtenbaum–Quillen and telescope questions in the case of algebraic K-theory of truncated Brown–Peterson spectra. Let $T(n)=\mathbb{S}/(p^{i_{0}},\cdots,v_{n-1}^{i_{n-1}})[v_{n}^{-1}]$ for suitably large integers $i_{0},\cdots,i_{n-1}$. We write $L_{n}$ for the Bousfield localization at $\mathbb{Q}\oplus K(1)\oplus\cdots K(n)$ where $K(i)$ is the $i$-th Morava K-theory at the prime $p$. We also write $L_{n}^{f}$ for the Bousfield localization at $\mathbb{Q}\oplus T(1)\oplus\cdots\oplus T(n)$.
Theorem A (Corollary 6.5, Corollary 6.6 and Corollary 6.7).
Let $p$ be a prime number and $n\geq 1$ an integer. Let $\operatorname{BP}\langle n\rangle$ be an arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of the $n$-th truncated Brown–Peterson spectra at the prime $p$, see Definition 2.1. Then
  1. (1)
    (Redshift) The spectrum $T(n+1)\otimes\operatorname{K}(\operatorname{BP}\langle n\rangle)$ is nonzero.
  2. (2)
    (Telescope) The localization map $L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)\to L_{n+1}\operatorname{K}(\operatorname{BP}\langle n\rangle)$ is an equivalence.
  3. (3)
    (Lichtenbaum–Quillen) The localization map $\operatorname{K}(\operatorname{BP}\langle n\rangle)\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)$ has bounded above fiber.
Each of these results follow from the following explicit computation.
Theorem B (Theorem 5.5 and Corollary 5.7).
Let $n\geq 1$ be an integer. The mod $(p,v_{1},\cdots,v_{n+1})$ syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
$\operatorname{\mathbb{F}}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle\oplus\bigoplus_{i=1}^{n+1}\operatorname{\mathbb{F}}_{p}\langle\lambda_{s}:s\in\langle n\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
as a bigraded $\operatorname{\mathbb{F}}_{p}$-vector space and the $v_{n+1}$-Bockstein spectral sequence collapses. Here the classes are in bidegrees $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ and $\langle n\rangle=\{1,2,\cdots,n\}$.
Here and throughout we write $\operatorname{\mathbb{F}}_{p}\langle x_{1},x_{2},\cdots,x_{n}\rangle$ for an exterior algebra with generators $x_{1},x_{2},\cdots,x_{n}$.
[…]
Figure 1. The mod $(2,v_{1},v_{2},v_{3})$-syntomic cohomology of $\operatorname{BP}\langle 2\rangle$
When $n=2$ and $p\geq 5$, we argue that the the motivic spectral sequence collapses and we can prove the following.
Theorem C (Theorem 6.1 and Corollary 6.2).
Let $p\geq 5$. We can identify
$\operatorname{TC}_{*}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})=\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\partial,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{i=1}^{3}\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\lambda_{s}:s\in\langle 3\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
where $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{j}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for each $1\leq j\leq 3$ and $0<d<p$ and $\langle 3\rangle=\{1,2,3\}$. There is a long exact sequence
$0\to\Sigma^{-2}\operatorname{\mathbb{F}}_{p}\{\tau_{1},\tau_{2},\tau_{1}\tau_{2}\}\to\operatorname{K}_{*}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})\to\operatorname{TC}_{*}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})\to\Sigma^{-1}\operatorname{\mathbb{F}}_{p}\{\partial\}\to 0$
where $|\tau_{i}|=2p^{i}-1$ for $i=1,2$.
Remark 1.1.
When n=1,0,1n=-1,0,1, Theorem A was first proven by [QUI72], [BM94], and [AR02] at the primes $p\geq 2$, $p\geq 3$, and $p\geq 5$ respectively. For arbitrary $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra Hahn and Wilson proved Theorem A in [HW22]. Our theorem extends this result to arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra, such as those arising from the choice of Araki or Hazewinkle generators.
When $n=0,1$, Theorem B was proven by [LW22] and [HRW22] respectively. As a consequence, Hahn–Raksit–Wilson extend Theorem A and the $n=1$ analogue of Theorem C to the prime $p=3$ in [HRW22].
When $n=2$, Theorem C was first proven by [AAC+25] at primes $p\geq 7$ for arbitrary $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms. We extend this result to the prime $p=5$ and to arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms.
When $n>2$, the answer for syntomic cohomology is exactly the answer predicted by Ausoni and Rognes [AR02, p. 5], as stated precisely in [ROG01, § 2], for topological cyclic homology of $\operatorname{BP}\langle n\rangle$ modulo $(p,v_{1},\cdots,v_{n})$. For their prediction to hold on the nose, one would need to show that the motivic spectral sequence collapses at the $\mathrm{E}_{2}$-term, which is not immediate.
Remark 1.2.
In fact, using Theorem B, T. Yang [YAN25, Theorem 2.7] has provided a precise bound in the statement of Theorem A (3), in particular the map
$\pi_{*}\operatorname{K}(\operatorname{BP}\langle n\rangle)\longrightarrow\pi_{*}L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)$
is an isomorphism in positive degrees.
Remark 1.3.
The syntomic cohomology computed in Theorem B and depicted in Figure 1 at $n=p=2$ suggests a certain rotational symmetry akin to the Lagrangian refinement of Tate duality of [BHA23]. A higher height analogue of this duality is the subject of ongoing work of Devalapurkar–Hahn–Rognes, see [ROG25].
1.1. Acknowledgements
The author is indebted to Jeremy Hahn and Dylan Wilson who contributed significantly to this project, but declined to be co-authors. Additionally, the author benefited from input from Christian Ausoni, Robert Burklund, Ishan Levy, Piotr Pstrągowski, Arpon Raksit, John Rognes, Andrew Senger, and Tristan Yang, either from direct communication or indirect communication though Jeremy Hahn and Dylan Wilson.
2. Hochschild homology
First, we define forms of $\operatorname{BP}\langle n\rangle$ in Section 2.1 and the $\operatorname{MU}$-based motivic filtration in Section 2.2. We compute Hochschild homology of $\operatorname{MU}$, $\operatorname{BP}$, and $\operatorname{BP}\langle n\rangle$ with $\operatorname{\mathbb{F}}_{p}$-coefficients in Section 2.3 along with their motivic filtrations in Section 2.4.
2.1. Forms of $\operatorname{BP}\langle n\rangle$
By [CM15, Corollary 1.3], the Quillen idempotent defines an $\operatorname{\mathbb{E}}_{2}$ algebra map
$\operatorname{BP}\to\operatorname{MU}_{(p)}\,.$
While $\pi_{*}\operatorname{BP}$ is only non-canonically isomorphic to a polynomial ring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots]$, for each $i\geq 1$ the subring $\mathbb{Z}_{(p)}[v_{1},\cdots,v_{i}]\subset\pi_{*}\operatorname{BP}$ is well-defined. Indeed, it is the subring generated by all elements of degree at most $2p^{i}-2$. Based on this, the following definition is standard:
Definition 2.1.
An $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle n\rangle$ is a $p$-local $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra $R$ such that the composite
$\mathbb{Z}_{(p)}[v_{1},\cdots,v_{n}]\subset\operatorname{BP}_{*}\subset\pi_{*}\operatorname{MU}_{(p)}\to\pi_{*}R$
is an isomorphism. The last map in this composite is the $p$-localized unit of the $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra structure.
In particular, the Eilenberg–MacLane spectrum $\mathbb{Z}_{(p)}$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 0\rangle$ and the Adams summand $\ell$ of $p$-local topological K-theory $\mathrm{ku}$ is a $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 1\rangle$.
In this paper, we will be interested in the algebraic $K$-theories of $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$. Of course, algebraic $K$-theory depends only on underlying $\operatorname{\mathbb{E}}_{1}$ $\mathbb{S}$-algebra structure, but we will exploit $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra structure when making our computations.
In the remainder of this paper, we fix particular (but arbitrary) $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$, according to the convention below. Since the choices are arbitrary, the theorems we prove about $\operatorname{BP}\langle n\rangle$ hold for all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms.
Convention 2.2.
For the remainder of this paper, we use the symbol $\operatorname{\mathrm{BP}\langle n\rangle}$ to denote a fixed (but arbitrary) $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{\mathrm{BP}\langle n\rangle}$.
We also fix for each $i\geq 1$ an indecomposable polynomial generator
$v_{i}\in\pi_{2p^{i}-2}\operatorname{BP}\,$,
and denote also by $v_{i}$ the image of this class under the map $\pi_{2p^{i}-2}\operatorname{BP}\to\pi_{2p^{i}-2}\operatorname{MU}$. We make these choices such that, for each $i>n$, the class $v_{i}$ maps to zero in $\pi_{2p^{i}-2}\operatorname{BP}\langle n\rangle.$
2.2. The motivic filtration
We review the definition of the motivic filtration relative to $\operatorname{MU}$, defined on the $\operatorname{THH}$ of any $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra by [HRW22].
Recollection 2.3.
Suppose that $R$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra. Then we follow [HRW22, Definition 4.2.1] in defining
$\mathrm{fil}_{\textup{mot}/\operatorname{MU}}^{*}\operatorname{THH}(R):=\mathrm{fil}_{\textup{ev}/\operatorname{THH}(\operatorname{MU})}^{*}\operatorname{THH}(R)\,$,
which is a filtered $\operatorname{\mathbb{E}}_{0}$-$\mathrm{gr}^{*}_{\textup{ev}}\operatorname{THH}(\operatorname{MU})$-algebra. Concretely,
$\mathrm{fil}^{*}_{\textup{mot}/\operatorname{MU}}\operatorname{THH}(R)=\mathrm{Tot}\left(\tau_{\geq 2*}\left({\operatorname{THH}(R)}\otimes_{\operatorname{THH}(\operatorname{MU})}{\operatorname{MU}}^{\otimes_{\operatorname{THH}(\operatorname{MU})}\bullet+1}\right)\right)\,.$
In this particular paper, we will always study motivic filtrations relative to $\operatorname{MU}$. Thus, we make the following further simplifying convention:
Convention 2.4.
Given a $\operatorname{THH}(\operatorname{MU})$-module $M$, we abbreviate $\mathrm{fil}^{*}_{\textup{ev}/\operatorname{THH}(\operatorname{MU})}M$ as $\mathrm{fil}^{*}_{\textup{mot}}M$ throughout. Therefore, if $R$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra, we write $\mathrm{fil}^{*}_{\textup{mot}}\operatorname{THH}(R)$ for $\mathrm{fil}^{*}_{\textup{mot}/\operatorname{MU}}\operatorname{THH}(R)$.
Warning 2.5.
For a general $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra $R$, using $\mathrm{fil}_{\textup{mot}}^{*}\operatorname{THH}(R)$ to denote $\mathrm{fil}^{*}_{\textup{mot}/\operatorname{MU}}\operatorname{THH}(R)$ may be in contradiction with the notations of [HRW22, PST23]. However, for the particular $R$ we study in this paper, all potential meanings of $\mathrm{fil}_{\textup{mot}}^{*}\operatorname{THH}(R)$ are equivalent.
Definition 2.6.
For each $i>0$, we let $\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}/v_{i}$ denote the cofiber of $v_{i}$ considered as a self-map of $\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}$. For $M$ a $\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}$-module, we then define
$M/v_{i}:=M\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}(\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}/v_{i}).$
We sometimes write $M/(p,\dots,v_{i})$ for the iterated tensor product
$M\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}(\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}/p)\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}\cdots\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}(\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}/v_{i}).$
Remark 2.7.
Via the natural map $\operatorname{MU}\to\operatorname{THH}(\operatorname{MU})$, any $\mathrm{gr}^{*}_{\textup{ev}}\operatorname{THH}(\operatorname{MU})$-module is also a $\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}$-module. Thus, if $R$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra, we may refer to
$\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(R)/(p,\cdots,v_{i}).$
Remark 2.8.
For each $i\geq 1$, our convention that $v_{i}\in\pi_{2p^{i}-2}\operatorname{MU}$ is in the image of the unit map
$\pi_{2p^{i}-2}\operatorname{BP}\to\pi_{2p^{i}-2}\operatorname{MU}$
ensures that the natural map
$\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}/(p,v_{1},\cdots,v_{i-1})\to\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}/(p,v_{1},\cdots,v_{i-1})$
sends $v_{i}$ to $v_{i}$. In particular, if $M$ is a $\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}$-module, then
$M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}}(\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}/(p,\cdots,v_{i}))\cong M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}}(\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}/(p,\cdots,v_{i})).$
Definition 2.9.
Given a graded spectrum $M^{*}$, we say $x\in\pi_{n}M^{w}$ has degree $n$, Adams weight 2wn2w-n and weight $w$. We write $\|x\|=(n,2w-n)$ and simply $|x|=n$ in this case.
2.3. Hochschild homology with $\operatorname{\mathbb{F}}_{p}$-coefficients
We begin with a non-motivic result, which is known for $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$, but is new for arbitrary $\operatorname{\mathbb{E}}_{1}$-$\operatorname{MU}$-algebra forms. First, recall the known computations of $\operatorname{THH}_{*}(\operatorname{MU};\mathbb{F}_{p})$ and $\operatorname{THH}_{*}(\operatorname{BP};\mathbb{F}_{p})$. Here and throughout given an $\operatorname{\mathbb{E}}_{1}$-ring $R$ and an $R\otimes R^{\textup{op}}$-module $M$ we write $\operatorname{THH}(R;M):=M\otimes_{R\otimes R^{\textup{op}}}R$.
Recollection 2.10.
There is an isomorphism
$\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})=\mathbb{F}_{p}[\mu]$
of graded $\operatorname{\mathbb{F}}_{p}$-algebras by [BÖK87], where $|\mu|=2$, and an isomorphism
$\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})=\mathbb{F}_{p}\langle\lambda_{i}^{\prime}:i\geq 1\rangle$
of graded $\operatorname{\mathbb{F}}_{p}$-algebras, where $|\lambda_{i}^{\prime}|=2i+1$ by [MS93, Remark 4.3]. We write $\lambda_{j}\coloneqq\lambda_{p^{j}-1}^{\prime}$. The $\operatorname{\mathbb{E}}_{3}$ algebra $\operatorname{THH}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})$ is a retract of $\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$ as an $\mathbb{E}_{1}$ algebra by [CM15, Corollary 1.3] and has homotopy groups
$\operatorname{THH}_{*}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})\cong\mathbb{F}_{p}\langle\lambda_{i}:i\geq 1\rangle\,.$
Property 2.11.
As an $\operatorname{\mathbb{E}}_{0}$ $\operatorname{THH}(\operatorname{BP})$-algebra,
$\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\cong\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}].$
Here, the class $\mu^{p^{n+1}}$ denotes a polynomial generator that maps to the similarly named class in $\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})$.
Proof.
Consider the Hochschild–May spectral sequence [AS18, KEE25, LL26] with $\mathrm{E}_{1}$-page
$\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p}[v_{0},\cdots,v_{n}];\operatorname{\mathbb{F}}_{p})$
as an $\operatorname{\mathbb{E}}_{0}$ algebra over the Hochschild–May spectral sequence for $\operatorname{BP}$ considered in  [AHW24b, Proposition 2.2.2]. Explicitly, we know $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{BP}$-algebra by Chadwick–Mandell [CM15, Corollary 1.3, Theorem 1.2] and consequently, we know that the filtered spectrum $B_{*}=\lim_{\Delta}\tau_{\geq 2*}(\operatorname{BP}\langle n\rangle\otimes\operatorname{\mathbb{F}}_{p}^{\otimes\bullet+1})$ is an $\operatorname{\mathbb{E}}_{1}$ $R_{*}$-algebra by Patchkoria–Pstrągowski [PP21] where $R_{*}=\lim_{\Delta}\tau_{\geq 2*}(\operatorname{BP}\otimes\operatorname{\mathbb{F}}_{p}^{\otimes\bullet+1})$. The $\mathrm{E}_{1}$-page can be identified with $\operatorname{\mathbb{F}}_{p}\langle\sigma v_{0},\cdots,\sigma v_{n}\rangle[\mu]$ by [HW18, Lemma 4.1.3] as an $\operatorname{\mathbb{E}}_{0}$ algebra over the $\mathrm{E}_{1}$-page of the Hochschild–May spectral sequence for $\operatorname{BP}$. Here the $\operatorname{\mathbb{E}}_{0}$ algebra structure is given by the canonical quotient map by $\sigma v_{j}$ for $j\geq n+1$. The $\operatorname{\mathbb{E}}_{0}$ algebra structure implies the differentials $d_{2p^{i}-2}(\mu^{p^{i}})=\sigma v_{i}$ and $d_{2p^{j}-2}(\mu^{p^{j}})=0$ as well as the differentials generated by the module structure. After considering these differentials as well as bidegrees, we observe that all remaining classes are infinite cycles. Again, we observe that there is no room for hidden $\operatorname{\mathbb{E}}_{0}$ $\operatorname{THH}(\operatorname{BP})$-algebra extensions. The map of Hochschild–May spectral sequences induced by $\operatorname{BP}\langle n\rangle\to\operatorname{\mathbb{F}}_{p}$ is the canonical quotient by $\sigma v_{i}$ for $0\leq i\leq n$ on the $\mathrm{E}_{1}$-page and we observe that it is the canonical quotient by $\lambda_{i}$ for $1\leq i\leq n+1$ followed by the canonical inclusion $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]\subset\operatorname{\mathbb{F}}_{p}[\mu]$ and therefore we name classes accordingly. ∎
This extends [ACH24a, Proposition 2.9] to arbitrary $\mathbb{E}_{1}$-$\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$.
2.4. The motivic filtration on Hochschild homology
The purpose of this section is to compute the homotopy groups of $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$, defined using the $\operatorname{THH}(\operatorname{MU})$-module structure on $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$. These bigraded homotopy groups form the $\mathrm{E}_{2}$-page of motivic spectral sequence
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\implies\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$
converging to the results of the previous subsection. We will deduce that the motivic spectral sequence degenerates at the $\mathrm{E}_{2}$-page without extensions.
Recall from [AHW24b, Proposition 2.2.6, Proposition 2.2.9] that
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{MU})=\operatorname{\mathbb{F}}_{p}\langle\lambda_{i}^{\prime}:i\geq 1\rangle$
and
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})=\operatorname{\mathbb{F}}_{p}\langle\lambda_{j}:j\geq 1\rangle$
where $\|\lambda_{i}^{\prime}\|=(2i-1,1)$, $\|\lambda_{j}\|=(2p^{j}-1,1)$ and using the Quillen idempotent we observe that the $\operatorname{MU}$-algebra structure on $\operatorname{BP}$ induces a map
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{MU})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})$
sending $\lambda_{i}^{\prime}$ to $\lambda_{j}$ when $i=p^{j}$ and 0 otherwise. Moreover, by [AHW24b, Proposition 2.2.8],
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})=\operatorname{\mathbb{F}}_{p}[\mu]$
where $\|\mu\|=(2,0)$.
First, we need a preliminary result, which uses the notion of even flatness due to Pstrągowski, see [PST23, Definition 4.4].
Property 2.12.
The $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-module $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is an even flat. In particular, the spectrum $\operatorname{THH}(\operatorname{BP}\langle n\rangle/\operatorname{MU})$ is even.
Proof.
It suffices to check that
$\operatorname{Tor}_{*}^{\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})}(\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p}),\operatorname{\mathbb{F}}_{p})$
is concentrated in even total degrees by [PST23, Proposition 4.20,Theorem 4.21]. This follows from Proposition 2.11. Consequently, we conclude that $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle/\operatorname{MU};\operatorname{\mathbb{F}}_{p})$ is concentrated in even degrees and since $|v_{i}|=2p^{i}-2$ the $v_{i}$-Bockstein spectral sequences for $i=0,1,\cdots,n$ are also concentrated in even degrees. Therefore $\operatorname{THH}(\operatorname{BP}\langle n\rangle/\operatorname{MU})$ is even. ∎
Proposition 2.13.
We can identify
$\pi_{*}\mathrm{gr}^{\star}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]$
as an $\operatorname{\mathbb{E}}_{0}$ $\pi_{*}\mathrm{gr}^{\star}_{\textup{mot}}\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebra where the bidegrees of generators are
$\|\mu^{sp^{n}}\|$ $=(2p^{n+1}s,0)$
$\|\lambda_{j}\|$ $=(2p^{j}-1,1)$
for $s\geq 0$ and $1\leq j\leq n+1$ and we write $\mu^{k}$ for a class that has image $\mu^{k}$ in
$\pi_{*}\mathrm{gr}^{\star}_{\textup{mot}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})=\operatorname{\mathbb{F}}_{p}[\mu]\,.$
The $\operatorname{\mathbb{E}}_{0}$ algebra structure map $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is given by
$\lambda_{i}^{\prime}\mapsto\begin{cases}\lambda_{j}&\text{ if }i=p^{j},j=1,\cdots,n+1\\ 0&\text{ otherwise}\end{cases}$
Proof.
By [AHW24b, Lemma 2.2.17] and Proposition 2.12, it suffices to observe that
$\operatorname{\operatorname{Cotor}}_{(\operatorname{\mathbb{F}}_{p},\Gamma)}^{*,*}(\operatorname{\mathbb{F}}_{p},\operatorname{THH}_{*}(\mathrm{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]$
as a $\operatorname{\mathbb{E}}_{0}$ $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebra and this identification is compatible with the map
$\operatorname{\operatorname{Cotor}}_{(A,\Gamma)}^{*,*}(A,\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}(\mathrm{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})\to\operatorname{\operatorname{Cotor}}_{(A,\Gamma)}^{*,*}(A,\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}(\mathrm{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})$
of $\operatorname{\mathbb{E}}_{0}$-$\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebras where
$\operatorname{\operatorname{Cotor}}_{(A,\Gamma)}^{*,*}(A,\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}(\mathrm{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})\cong\operatorname{\mathbb{F}}_{p}[\mu]\,.$
Proposition 2.14.
There is an equivalence of $\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{MU})$-modules
$\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\simeq\left(\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)\right)/(p,v_{1},\cdots,v_{n})$
Proof.
By Proposition 2.12, we know that $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is even. Therefore, the $p,v_{1},\cdots,v_{n}$-Bockstein spectral sequences all collapse and we conclude that $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle/\operatorname{MU})_{p}^{\wedge}$ is a finitely generated free $\operatorname{\mathbb{Z}}_{p}^{\wedge}[v_{1},\cdots v_{n}]$-module. Letting $B_{i}=\operatorname{BP}\langle i\rangle$ for $1\leq i\leq n$. We have cofiber sequences
$\mathrm{gr}_{\textup{mot}}^{*}(\Sigma^{2p^{i}-2}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i-1}))\to\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i-1}))\to\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i})$
induced by the short exact sequences
$0\to\pi_{*}\Sigma^{2p^{i}-2}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i-1})\to\pi_{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i-1})\to\pi_{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i})\to 0\,.$
We can identify
$\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i})\simeq\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;B_{i-1})/(v_{i})$
for each $0\leq i\leq n$ and after a finite induction we have proven the claim. ∎
Theorem 2.15.
There is an isomorphism
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\cong\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]\langle\varepsilon_{n+1}\rangle$
of $\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle$-modules and the canonical map
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n+1})$
of $S^{1}$-equivariant $\operatorname{\mathbb{E}}_{0}$ $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{MU})/(p,v_{1},\cdots,v_{n})$-algebras is injective modulo $(\lambda_{1},\cdots,\lambda_{n+1})$ with image exactly $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n+1}}]\langle\varepsilon_{n+1}\rangle$. Here the bidegree of $\varepsilon_{n+1}$ is
$\|\varepsilon_{n+1}\|=(2p^{n+1}-1,-1)\,.$
Proof.
The canonical map
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n})$
factors through the map
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$
with image $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]$ by Proposition 2.14 and Proposition 2.11. This determines the image of the map
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n})$
as well. Taking a further quotient by $v_{n+1}$ produces the result. To see this, consider the commutative diagram of $\operatorname{THH}(\operatorname{MU})$-modules
$\operatorname{MU}$$\operatorname{THH}(\operatorname{MU})$$\operatorname{BP}\langle n\rangle$$\operatorname{THH}(\operatorname{BP}\langle n\rangle)$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$}
to determine that the $v_{n+1}\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}$ acts trivially on $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)$ compatibly with the map $\operatorname{THH}(\operatorname{BP}\langle n\rangle)\to\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$. ∎
3. Hodge Tate Cohomology
The main theorem of this section is the following calculation of mod $(p,v_{1},\cdots,v_{n+1})$ Hodge–Tate cohomology:
Theorem 3.1.
On bigraded homotopy groups, the commutative diagram
$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})\,$$\scriptstyle{\varphi_{p}}$$\scriptstyle{\varphi_{p}}$
where the vertical maps are induced by the reduction map $\operatorname{BP}\langle n\rangle\to\operatorname{\mathbb{F}}_{p}$ and the horizontal maps are the Frobenius maps, may be identified as $\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle$-modules with the square
$\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]\langle\varepsilon_{n+1}\rangle$$\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle[\mu^{\pm p^{n+1}}]\langle\varepsilon_{n+1}\rangle$$\operatorname{\mathbb{F}}_{p}[\mu]\langle\epsilon_{0},\epsilon_{1},\cdots,\epsilon_{n+1}\rangle$$\operatorname{\mathbb{F}}_{p}[\mu^{\pm 1}]\langle\epsilon_{0},\epsilon_{1},\cdots,\epsilon_{n+1}\rangle$}
where the horizontal maps are given by inverting $\mu^{p^{n+1}}$ and $\mu$ respectively and the vertrical maps are the tensor product of the canonical quotient $\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle\to\operatorname{\mathbb{F}}_{p}$ with the canonical inclusion.
Proof.
First, note that there is a canonical map
$L_{n}^{f}\operatorname{MU}\longrightarrow L_{n}^{f}\operatorname{BP}\langle n\rangle$
and since $\tau_{\geq 0}L_{n}^{f}\operatorname{BP}\langle n\rangle\simeq\operatorname{BP}\langle n\rangle$ there is a canonical map
$\tau_{\geq 0}L_{n}^{f}\operatorname{MU}\longrightarrow\tau_{\geq 0}L_{n}^{f}\operatorname{BP}\langle n\rangle\simeq\operatorname{BP}\langle n\rangle\,.$
Since $v_{n+1}\in\pi_{2p^{n+1}-2}\operatorname{MU}$ maps trivially in $\pi_{2p^{n+1}-2}L_{n}^{f}\operatorname{MU}$, we can consider the element
$\sigma^{2}v_{n+1}\in\pi_{2p^{n+1}}\operatorname{THH}(\tau_{\geq 0}L_{n}^{f}\operatorname{MU}/\operatorname{MU})$
which is non-zero because it maps to a non-zero class
$\sigma^{2}v_{n+1}=\mu^{p^{n+1}}\in\pi_{2p^{n+1}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p}/\operatorname{MU})\,.$
This produces a element $\mu^{p^{n+1}}$ in the bigraded commutative commutative ring
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\tau_{\geq 0}L_{n}^{f}\operatorname{MU})$
which acts on $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)$. It therefore makes sense to consider the map
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\longrightarrow\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p}^{\otimes p})^{tC_{p}}$
as a map of $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]$-algebras and it makes sense to ask whether the map is given by inverting the class $\mu^{p^{n}}$. As in [HW22, Recollection 4.3.2], we can determine that the associated graded $\mathrm{gr}_{\textup{Ad}}^{*}\operatorname{BP}\langle n\rangle$ of the filtration
$\lim_{\Delta}\tau_{\geq*}(\operatorname{BP}\langle n\rangle\otimes\operatorname{\mathbb{F}}_{p}^{\otimes\bullet+1})$
can be identified with $\operatorname{\mathbb{F}}_{p}\otimes\operatorname{\mathbb{S}}[v_{0},\cdots,v_{n}]$ as $\operatorname{\mathbb{E}}_{1}$-algebras. We can apply the same construction to $\tau_{\geq 0}L_{n}^{f}\operatorname{MU}$ and produce an action of $\mu$ on the map of $\mathrm{E}_{2}$-pages
$\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{\mathbb{F}}_{p}}\pi_{*}\operatorname{\mathbb{F}}_{p}\otimes\operatorname{THH}(\operatorname{\mathbb{S}}[v_{0},\dots,v_{n}];\operatorname{\mathbb{S}})$$\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})^{tC_{p}}\otimes_{\operatorname{\mathbb{F}}_{p}}\pi_{*}\operatorname{\mathbb{F}}_{p}\otimes\operatorname{THH}(\operatorname{\mathbb{S}}[v_{0},\dots,v_{n}];\operatorname{\mathbb{S}}^{\otimes p})^{tC_{p}}$}
and in fact the same considerations as in the proof of [HW22, Theorem 4.3.1] allow us to determine that the map is given by inverting $\mu$ on the $\operatorname{\mathbb{E}}_{2}$-pages. We have completely computed the top spectral sequence and since the bottom spectral sequence is given by inverting $\mu$ on $\operatorname{\mathbb{E}}_{2}$-pages, we determine that at each page the spectral sequence is given by inverting some power of $\mu$, and hence remains a monomomorphism entirely determining also the spectral sequence on the bottom. This proves the claim. ∎
Theorem 3.2.
The spectral sequence beginning with
$\left(\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})\right)[t]$
and converging to $\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$ collapses after the $d_{1}$ differential. All permanent cycles are on the 0-line.
Proof.
Since $v_{n+1}$ is detected by $t\mu^{p^{n+1}}$, we determine that $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$. Using the action of $\lambda_{1},\cdots,\lambda_{n+1}$ together with the monomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\cdots,\lambda_{n+1})$, we determine a Leibniz rule $d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$ for all
$x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n})\,.$
Since $\mu$ is a unit in a ring that acts on this spectral sequence, mainly the corresponding spectral sequence for $\tau_{\geq 0}L_{n}^{f}\operatorname{MU}$, the spectral sequence collapses after running the first differential and all permanent cycles are concentrated on the 0-line. Moreover, we can conclude that
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)\cong\ker(\sigma)$
where
$\sigma:\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}\to\pi_{*+1}\mathrm{gr}_{\textup{mot}}^{*+1}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}$
is the $\sigma$-operator defined as in [AHW24b, § 2.1.2] using the $\mathrm{fil}_{\textup{ev}}^{*}\mathbb{S}[S^{1}]$-action. ∎
4. Prismatic cohomology
Recall that there is a map
$\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n+1})$
of $\operatorname{\mathbb{E}}_{0}$-$\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})$-algebras induced by the canonical $\operatorname{\mathbb{E}}_{1}$-$\operatorname{BP}$-algebra reduction map
$\operatorname{BP}\langle n\rangle\to\tau_{\leq 0}\operatorname{BP}\langle n\rangle=\operatorname{\mathbb{Z}}_{(p)}\to\operatorname{\mathbb{F}}_{p}.$
We computed that this map is injective mod $(\lambda_{1},\cdots,\lambda_{n+1})$, with image $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]\otimes\Lambda(\varepsilon_{n+1})$, in Theorem 2.15. This is the key fact needed to produce the following result. First, we need some notation.
Notation 4.1.
We write
$\lambda_{W}=\prod_{w\in W}\lambda_{w}$
for $W\subset\{1,\cdots,n+1\}$ where $\lambda_{W}=1$ if $W=\emptyset$.
Proposition 4.2.
The periodic $t$-Bockstein spectral sequence
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})[t^{\pm 1}]\otimes\Lambda(\varepsilon_{n+1})$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$}
has differentials
$d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}\,$,
for all $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})[t^{\pm 1}]$ and
$d_{p^{m}}(t^{p^{m-1}})=t^{p^{m}+p^{m-1}}\lambda_{m}$
for all $1\leq m\leq n+1$ as well as those differentials generated by the Leibniz rule
$d_{p^{m}}(t^{jp^{m-1}}\lambda_{S})\dot{=}t^{p^{m}+jp^{m-1}}\lambda_{m}\lambda_{S}$
for $\ell\in\mathbb{Z}$, $0<j<p$ and $S,T\subset\{1,\cdots,n+1\}$ with $m\not\in S$. The spectral sequence then collapses at the $p^{n+1}+1$-page without room for further differentials. Consequently,
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\cong\mathbb{F}_{p}\langle\lambda_{1},\dots,\lambda_{n+1}\rangle[t^{\pm p^{n+1}}]$
as $\mathbb{F}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle$-modules.
Proof.
The differential $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ follows from the definition of $\varepsilon_{n+1}$ and Theorem 2.15. The Leibniz rule
$d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$
for all $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})[t^{\pm 1}]$ follows from [AHW24b, Corollary 3.2.6], which implies that the classes $\lambda_{1},\cdots,\lambda_{n+1}$ are permanent cycles that act on this spectral sequence, together with the monomomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\cdots,\lambda_{n+1})$.
The differentials
$d_{p^{m}}(t^{p^{m-1}}\lambda_{S})=t^{p^{m}+p^{m-1}}\lambda_{m}\lambda_{S},$
for all $S\subset\{1,\cdots,n+1\}$ with $m\not\in S$ for each $m\geq 1$ follow from the action of the spectral sequence for $\operatorname{BP}$ and the differentials in that spectral sequence, computed in [AHW24b, Proposition 3.2.9]. Finally, we simply check bidegrees and observe that the spectral sequence must collapse at the $\mathrm{E}_{p^{n+1}+1}$-page. ∎
5. Syntomic cohomology
We now have all the ingredients to compute the syntomic cohomology of $\operatorname{BP}\langle n\rangle$. First, we need to determine
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\,.$
We note that these bigraded groups are equipped with the Nygaard filtration coming the $t$-Bockstein spectral sequence and therefore there is an exact sequence
$0\to\mathrm{Nyg}_{\geq 1}\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\mathrm{Nyg}_{=0}\to 0$
where $\mathrm{Nyg}_{\geq 1}$ denotes the subgroup generated by classes in positive Nygaard filtration and $\mathrm{Nyg}_{=0}$ is defined to be the cokernel. Let $T=\ker(\mathrm{can})$ and let $F=\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/\ker(\mathrm{can})$.
Notation 5.1.
There is a commutative diagram
[…]
where
$A_{10}$ $:=\mathrm{Nyg}_{\geq 1}\cap T\,$,
$A_{00}$ $:=\mathrm{Nyg}_{\geq 1}/A_{10}\,$,
$A_{11}$ $:=T/A_{10}\text{ and }$
$A_{10}$ $=F/A_{00}\,.$
Property 5.2.
The classes in the $\mathrm{E}_{2}$-page of the periodic $t$-Bockstein spectral sequence that are hit by differentials that cross from negative Nygaard filtration to positive Nygaard filtration are exactly
$A_{10}:=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,.$
Proof.
This follows from a careful bookkeeping of the differentials determined in Proposition 4.2. ∎
The following corollary is immediate.
Corollary 5.3.
We compute that
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{p^{n+1}}]\oplus$
$\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{d,j}:0<d<p\}$
as a $\mathbb{F}_{p}\langle\lambda_{1},\cdots\lambda_{n+1}\rangle$-module where $\Xi_{d,j}$ is detected by $t^{dp^{j-1}}\lambda_{j}$, $y^{p^{n+1}}$ is detected by $t^{-p^{n+1}}$ and we abuse notation and write $\lambda_{k}$ for the class detected by $\lambda_{k}$. Moreover, the canonical map
$\textup{can}:\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
sends $\lambda_{k}$ to $\lambda_{k}$, $y^{p^{n+1}}$ to $y^{p^{n+1}}$, and sends $\lambda_{s}^{\epsilon}\cdot\Xi_{d,j}$ to zero for each $\epsilon\in\{0,1\}$, $s\in\{1,\cdots,n\}-\{j\}$, and $j\in\{1,\cdots,n+1\}$.
We further need to compute the Frobenius map. We proceed by considering the map of $t$-Bockstein spectral sequences
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})[t]$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})[t]$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\,.$}
Property 5.4.
The $t$-Bockstein spectral sequence
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n+1})[t]\implies\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
has $\mathrm{E}_{2}$-term
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\otimes\mathbb{F}[\mu^{\pm p^{n}},t]\langle\varepsilon_{n+1}\rangle$
and it collapses after the differential
$d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$
along with those differentials given by the Leibniz rule
$d_{1}(\varepsilon_{n+1}x)=t\mu^{p^{n+1}}x$
where $x\in\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{\pm p^{n}},t]$. We can therefore identify
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{\pm p^{n+1}}]$
where $y^{p^{n+1}}$ is detected by $\mu^{p^{n+1}}$.
Letting
$A$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
$B$ $=\bigoplus_{j=1}^{n+1}\bigotimes_{s\in\{1,\cdots,n+1\}-\{j\}}\mathbb{F}_{p}\langle\lambda_{s}\rangle\{\Xi_{j,d}:0<d<p\}\,,\text{ and }$
$C$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{p^{n+1}}]\{y^{p^{n+1}}\}$
the Frobenius map
$\varphi:\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
is given by
$\varphi|_{A}=\textup{inc}_{A}\circ\mathrm{id}_{A}$
and $\varphi|_{B\oplus C}=0$ where $\textup{inc}_{A}:A\to A\oplus B\oplus C$ is the canonical inclusion.
We now present our computation of syntomic cohomology of $\operatorname{BP}\langle n\rangle$.
Theorem 5.5.
The mod $(p,\dots,v_{n+1})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
(5.6) $\mathbb{F}_{p}\langle\delta,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
as a $\mathbb{F}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module. In particular, it is a finite bigraded free $\operatorname{\mathbb{F}}_{p}$-module of dimension $2^{n+2}+2^{n}(n+1)(p-1)$ with generators concentrated in degrees $1,\sum_{i=1}^{n+1}2p^{i}-n-1]$ and Adams weights $[0,n+2]$. Here the classes are in bidegrees $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ for $1\leq i,j\leq n+1$ and $0<d<p$.
Proof.
Let
$A$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
$B$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}\,$,
$C$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{-p^{n+1}}]\{y^{-p^{n+1}}\}$
then we determined that $\mathrm{can}|_{A}=\textup{incl}_{A\oplus C}\circ\mathrm{id}_{A\oplus C}$ and $\mathrm{can}(B)=0$ in Corollary 5.3, where $\textup{incl}_{A\oplus C}$ denotes the canonical inclusion $A\oplus C\to A\oplus B\oplus C$. We determined that $\varphi|_{A}=\textup{inc}_{A}\circ\mathrm{id}_{A}$ and $\varphi|_{B\oplus C}=0$ in Proposition 5.4. We therefore determine that $\ker(\mathrm{can}-\varphi)=A\oplus B$ and $\operatorname{coker}(\mathrm{can}-\varphi)=A$. We write $\partial$ to indicate that these classes come from the boundary map in the long exact sequence mod $(p,v_{1},\cdots,v_{n+1})$ induced by the fiber sequence
$\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)\to\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)\overset{\mathrm{can}-\varphi}{\longrightarrow}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)\,.$
We present our answer using the notation from [AAC+25, Theorem 1.1]. ∎
Corollary 5.7.
The mod $(p,\dots,v_{n})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
(5.8) $\mathbb{F}_{p}[v_{n+1}]\langle\delta,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}[v_{n+1}]\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}.$
In particular, it is a bigraded free $\operatorname{\mathbb{F}}_{p}[v_{n+1}]$-module of dimension $2^{n+2}+2^{n}(n+1)(p-1)$ with generators concentrated in degrees $1,\sum_{i=1}^{n+1}2p^{i}-n-1]$ and Adams weights $[0,n+2]$. Here the classes are in bidegrees $\|v_{n+1}\|=(2p^{n+1}-2,0)$, $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ for $1\leq i,j\leq n+1$ and $0<d<p$.
Proof.
First, note that the $v_{n+1}$-Bockstein differentials raise Adams weight by exactly $1$ so it suffices to consider one Adams weight at a time. We then observe that $\mathbb{F}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$ is a subring of $\pi_{*}\operatorname{TC}(\operatorname{MU})/p$ by [AHW24b, Definition 3.2.4, Lemma 3.2.5.] so the differentials are $\partial,\lambda_{1},\cdots,\lambda_{n+1}$-linear. It therefore suffices to show that there are no $v_{n}$-Bockstein differentials on the generators of the $\mathrm{E}_{1}$-page of the $v_{n+1}$-Bockstein spectral sequence as a $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module.
The only generator in Adams weight 0 is $1$ and this cannot be the source of a $v_{n+1}$-Bockstein differential because the $\mathbb{F}_{p}[v_{n+1}]$-module generators of the $\mathrm{E}_{1}$-page of the $v_{n+1}$-Bockstein spectral sequence are concentrated in degrees $1,\sum_{i=1}^{n+1}2p^{i}-n-1]$ by Theorem 5.5.
The generators in Adams weight $1$ are $\Xi_{j,d}$ for $0<d<p$ and $1\leq j\leq n+1$. The classes in Adams weight $2$ are $\partial\lambda_{k}$ for $k\geq 1$, $\lambda_{i}\lambda_{j}$ for $i\neq j\in\{1,\cdots,n+1\}$, and $\lambda_{s}\Xi_{j,d}$ for $s\in\{1,\cdots,n+1\}-\{j\}$, $0<j<p$, and $0<d<p$. The class of lowest degree in Adams weight $2$ is $\partial\lambda_{1}$ in degree 2p22p-2 and the class in highest degree in Adams weight $1$ is $\Xi_{n+1,1}$ in $2p^{n+1}-3$, so since
$|\partial\lambda_{1}|+|v_{n+1}|+1=2p-2+2p^{n+1}-2+1>2p^{n+1}-3$
for all primes $p>2$. This handles all the $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module generators and therefore proves the claim and therefore the $v_{n+1}$, $\partial$, $\lambda_{1}$, $\cdots$ $\lambda_{n+1}$-linearity proves the claim. ∎
Remark 5.9.
The dimension of the syntomic cohomology of $\operatorname{BP}\langle n\rangle$ mod $(p,v_{1},\cdots,v_{n+1})$ was predicted by Ausoni–Rognes [AR02, p. 7], as stated precisely in [ROG01, §2]. They also claim that when $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{\infty}$-ring and a Smith–Toda complex $\operatorname{\mathbb{S}}/(p,v_{1},\cdots,v_{n})$ exists, then the $v_{n+1}$-Bockstein spectral sequence and the motivic spectral sequence collapse. It follows from [AAC+25], that at least when $p\geq 7$ and $n=2$ it suffices to know that $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{3}$-ring, which is known by [HW22].
6. Algebraic K-theory
In this section, we compute algebraic K-theory of $\operatorname{BP}\langle n\rangle$ at sufficiently low heights. We then prove Lichtenbaum—Quillen properties for algebraic K-theory of $\operatorname{BP}\langle n\rangle$. Finally, we prove that the telescope conjecture is true for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
6.1. Algebraic K-theory at low heights
We provide some immediate applications of our work to computations of topological cyclic homology of $\operatorname{BP}\langle 2\rangle$ at primes $p\geq 5$.
Theorem 6.1.
Let $p\geq 5$. Then $\pi_{*}\operatorname{TC}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})$ is isomorphic to
$\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}[v_{3}]\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
as $\mathbb{F}_{p}[v_{3}^{d}]$-modules for some integer $d\geq 1$. Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
Proof.
For simplicity, we display the mod $(p,v_{1},v_{2},v_{3})$-syntomic cohomology of $\operatorname{BP}\langle 2\rangle$ in Figure 1, but the reader should be able to extrapolate the necessary information from this picture to verify the statements in this proof. Since the motivic spectral sequence is concentrated on lines 0, 11, 22, 33 and $4$ the only possible motivic differential is one of length $3$ from 0-line to the $3$-line or the $1$-line to $4$-line. For differentials from the 0-line to the $3$-line, the only possible differentials have source $v_{3}^{k}$, but the targets are bidegree $((2p^{3}-2)k,3)$ and the mod $(p,v_{1},v_{2})$-syntomic cohomology of $\operatorname{BP}\langle 2\rangle$ is trivial in these bidegrees. Similarly, the only possible nontrivial targets of a differential of length $3$ from the $1$-line to the $4$-line are $\lambda_{1}\lambda_{2}\lambda_{3}v_{3}^{k}$, but there is a gap between $(2p^{3}+2p^{2}+2p-1,1)$ and $(2p^{3}+4p^{2}+2p-1,1)$ so there could not be any differential hitting $\delta\lambda_{1}\lambda_{2}\lambda_{3}$ and this gap persists in positive dimensions mod $2p^{3}-2$ that there are not possible differentials hitting $\lambda_{1}\lambda_{2}\lambda_{3}v_{3}^{k}$. Since we can view $\operatorname{TC}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})$ as a $\textup{End}(\mathbb{S}/(p,v_{1},v_{2}))$-module, which has a $v_{3}^{d}$-self map for some $d\geq 1$ and $\pi_{0}\textup{End}(\mathbb{S}/(p,v_{1},v_{2}))=\mathbb{F}_{p}$, this identification is an identification of $\mathbb{F}_{p}[v_{3}^{d}]$-modules ∎
Corollary 6.2.
Let $p\geq 5$. We compute
$\pi_{s}\mathrm{K}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})\cong\begin{cases}0&s<0\\ \pi_{s}\mathrm{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})\oplus\mathbb{F}_{p}\{\Sigma^{-2}\overline{\tau}_{i}\}&s=2p^{i}-3,i=1,2\\ \pi_{s}\mathrm{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})\oplus\mathbb{F}_{p}\{\Sigma^{-2}\overline{\tau}_{1}\overline{\tau}_{2}\}&s=2p^{2}+2p-4\\ \pi_{s}\mathrm{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})&\text{ otherwise }\end{cases}$
and $\pi_{*}\mathrm{K}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})[v_{3}^{-1}]$ is isomorphic to
$\operatorname{\mathbb{F}}_{p}[v_{3}^{\pm 1}]\otimes\bigl(\mathbb{F}_{p}\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}\bigr)$
as $\mathbb{F}_{p}[v_{3}^{\pm d}]$-modules for some integer $d\geq 1$. Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
Proof.
By [DGM13, Theorem 3.1.14], there is a fiber sequence
$\mathrm{K}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})\to\mathrm{TC}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})\to\Sigma^{-1}\mathbb{Z}/(p,v_{1},v_{2})$
where $\pi_{*}\mathbb{Z}/(p,v_{1},v_{2})=\mathbb{F}_{p}\langle\overline{\tau}_{1},\overline{\tau}_{2}\rangle$. This already implies the second claim. We know that the classes $\partial\lambda_{i}$ for $i=1,2$ and lift to mod $p$ homotopy classes so by the commuting diagram
$\mathrm{TC}(\operatorname{BP}\langle 2\rangle)/(p)$$\Sigma^{-1}\mathbb{Z}/p=\Sigma^{-1}\mathbb{F}_{p}$$\mathrm{TC}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})$$\Sigma^{-1}\mathbb{Z}/(p,v_{1},v_{2})$}
the map $\operatorname{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})\to\Sigma^{-1}\mathbb{Z}/(p,v_{1},v_{2})$ is the zero map in non-negative degrees. Since mod $(p,v_{1},v_{2})$-algebraic K-theory of $\operatorname{BP}\langle n\rangle$ is connective, this implies the first claim. ∎
Remark 6.3.
This extends work of the author with Ausoni, Culver, Höning, and Rognes [AAC+25, Theorem 1.2] to all $\operatorname{\mathbb{E}}_{1}$-$\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle 2\rangle$. It also extends the computation to the prime $p=5$.
6.2. Lichtenbaum–Quillen
In this section, we note that our computations imply various Lichtenbaum–Quillen type properties.
Theorem 6.4.
For any integers $i_{0},\cdots,i_{n+1}$ such that the generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$ exists, the graded abelian group
$\pi_{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
is finite; i.e. it is finite in each degree and only non-trivial in finitely many degrees. Consequently, the algebraic K-theory $\mathrm{K}(\mathrm{BP}\langle n\rangle)$ of $\mathrm{BP}\langle n\rangle$ has fp-type $n+1$.
Proof.
By Theorem 5.5, the bigraded abelian group
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p,\cdots,v_{n+1})$
is finite in the sense it is finite in all bidegrees and only non-trivial in finitely many bidegrees. This immediately implies that
$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
is finite for any integers $i_{0},\cdots,i_{n+1}$. Choosing integers $i_{0},\cdots,i_{n+1}$ such that a generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$ exists, this immediately implies that
$\pi_{*}\operatorname{TC}(R)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
is finite by motivic spectral sequence and consequently $\operatorname{TC}(R)$ has fp-type $n+1$ in the sense of [MR99]. By [DGM13, Theorem 3.1.14], there is a fiber sequence
$\operatorname{K}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})\to\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})\to\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
the algebraic K-theory of $R$ has fp-type $n+1$ since $\pi_{*}\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ is also finite. ∎
Remark 6.1.
This extends [HW22, Theorem B] to arbitrary $\mathbb{E}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$.
This implies the higher height analogue of the Lichtenbaum–Qullen conjecture as proposed by Ausoni and Rognes in [AR08].
Corollary 6.5.
The the fiber of the map
$\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}$
is bounded above.
Proof.
This follows from Theorem 6.4 by [MR99, Theorem 8.2(2)] as observed in [HW22, Theorem 3.1.3]. ∎
6.3. Telescope
We now note that our results also imply the telescope conjecture for topological cyclic homology of $\operatorname{BP}\langle n\rangle$.
Corollary 6.6.
The localization map
$L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)\to L_{n+1}\operatorname{K}(\operatorname{BP}\langle n\rangle)$
is an equivalence.
Proof.
Consider the $S^{1}$-equivariant eff map $\operatorname{THH}(\operatorname{MU})\to\operatorname{MU}$ by [HRW22, Example 4.2.3]. It is clear that $\pi_{*}\operatorname{BP}\langle n\rangle_{p}^{\wedge}$ is a finitely generated $\operatorname{\mathbb{Z}}_{p}$-module and that $R$ has height $n$ in either case. We computed that $\operatorname{BP}\langle n\rangle_{p}^{\wedge}$ has bounded below topological Hochschild homology in Proposition 2.11, the $\operatorname{THH}(\operatorname{MU})$-module $\operatorname{THH}(\operatorname{BP}\langle n\rangle)$ has bounded $\operatorname{MU}$-based motivic cohomological dimension in Proposition 2.13 and that $\operatorname{THH}(\operatorname{BP}\langle n\rangle/\operatorname{MU})=\operatorname{THH}(R)\otimes_{\operatorname{THH}(\operatorname{MU})}\operatorname{MU}$ is even in Proposition 2.12. Therefore, the result follows from [AHW24b, Proposition 4.3.2]. ∎
6.4. Redshift
We now prove redshift for algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
Corollary 6.7.
The algebraic K-theory of $\operatorname{BP}\langle n\rangle$ has height exactly $n+1$.
Proof.
By [LMM+24, Corollary 4.12], it suffices to prove that the algebraic K-theory of $\operatorname{BP}\langle n\rangle$ has height at least $n+1$. Since $\operatorname{\mathbb{E}}_{0}$ algebra structure on $\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)$ factors
$\operatorname{\mathbb{S}}\to\operatorname{K}(\operatorname{BP}\langle n\rangle)\to\operatorname{TC}(\operatorname{BP}\langle n\rangle)\to\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle/\operatorname{MU})$
it suffices to prove that a class in motivic filtration 0 of $\operatorname{TC}(\operatorname{BP}\langle n\rangle)$ in the kernel of $\mathrm{can}-\varphi$ is not $(p^{i_{0}},\cdots,v_{n}^{i_{n}})$-torsion for any postive integers $i_{0},\cdots,i_{n}$. and that this class is $v_{n+1}$-periodic. Choosing the class $1\in\pi_{0}\operatorname{TC}(\operatorname{BP}\langle n\rangle)$ in the image of the element $1\in\pi_{0}\operatorname{\mathbb{S}}$ then we can conclude that this class has both of these properties by Corollary 5.7. ∎

### v2

@@
 As qualitative consequences, we resolve the Lichtenbaum–Quillen, telescope, and redshift questions for the algebraic K-theories of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$.
-This extends work of the Hahn and Wilson.
-We also explicitly compute the algebraic K-theory of arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle 2\rangle$ at all primes $p\geq 5$ extending previous work of the author, Ausoni, Culver, Höning, and Rognes.
+This extends work of Hahn and Wilson.
+We also explicitly compute the mod $(p,v_{1},v_{2})$ algebraic K-theory of arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle 2\rangle$ at all primes $p\geq 5$ extending previous work of the author, Ausoni, Culver, Höning, and Rognes.
 1. Introduction
@@
 By Lazard [LAZ55], $p$-typical (1-dimensional) formal group laws are classified by their height and we may ask whether there is a complex oriented homology theory that carries a formal group law of height $\leq n$.
-One way to build such a ring spectrum is to take a quasi-regular quotient of Brown–Peterson homology to produce the truncated Brown-Peterson homology $\mathrm{BP}\langle n\rangle$ with coefficients a polynomial algebra over the $p$-local integers on generators $v_{1}$, $v_{2}$, $\dots$ $v_{n}$.
+One way to build such a ring spectrum is to take a quasi-regular quotient of Brown–Peterson homology to produce truncated Brown-Peterson homology $\mathrm{BP}\langle n\rangle$ with coefficients a polynomial algebra over the $p$-local integers on generators $v_{1}$, $v_{2}$, $\dots$ , $v_{n}$.
+By convention $\mathrm{BP}\langle-1\rangle=\mathbb{F}_{p}$.
 Structured ring spectra play a fundamental role in homotopy theory, for example more highly structured ring spectra are equipped with a richer notion of power operations.
@@
 It was hoped that the Brown–Peterson homology theory and the truncated Brown–Peterson cohomology theories would have similar structure.
-It was shown that $\operatorname{\mathbb{E}}_{\infty}$ algebra structures exists on $\operatorname{BP}\langle n\rangle$ for n=1,0,1n=-1,0,1 at all primes and $n=2$ at the primes $p=2,3$ by [BR08, HL10, LN12, LN14] and an $\operatorname{\mathbb{E}}_{4}$-algebra structure exists on $\operatorname{BP}$ by [BM13].
-However, it was proven by [LAW18, SEN24] that $\operatorname{BP}$ and $\operatorname{BP}\langle n\rangle$ do not have $\operatorname{\mathbb{E}}_{\infty}$ algebra structures when $n\geq 4$, so we do not live in the best of all possible worlds.
-Producing $\operatorname{\mathbb{E}}_{m}$ algebra structures on truncated Brown–Peterson spectra for $m>1$ at arbitrary heights $n$ remained an open problem until [HW22] proved that there exist $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra for all primes $p$ and heights $n$.
+It was shown that $\operatorname{\mathbb{E}}_{\infty}$-ring structures exist on $\operatorname{BP}\langle n\rangle$ for n=1,0,1n=-1,0,1 at all primes and $n=2$ at the primes $p=2,3$ by [BR08, HL10, LN12, LN14] and an $\operatorname{\mathbb{E}}_{4}$-ring structure exists on $\operatorname{BP}$ by [BM13].
+However, it was proven by Lawson and Senger [LAW18, SEN24] that $\operatorname{BP}$ and $\operatorname{BP}\langle n\rangle$ do not have $\operatorname{\mathbb{E}}_{\infty}$-ring structures when $n\geq 4$, so we do not live in the best of all possible worlds.
+Producing $\operatorname{\mathbb{E}}_{m}$-ring structures on truncated Brown–Peterson spectra for $m>1$ at arbitrary heights $n$ remained an open problem until Hahn and Wilson [HW22] proved that there exist $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra for all primes $p$ and heights $n$.
 These forms depend on a choice of generators $v_{i}$.
 The most common choice would be to choose the Araki or Hazewinkle generators.
-However, it was observed by Strickland [STR99, Remark 6.5] at the prime $2$ that, although these forms of truncated Brown–Peterson spectra are equipped with $\operatorname{\mathbb{E}}_{1}$ algebra structures, they are not equipped with additional structure.
-In particular, they are not equipped with $\operatorname{\mathbb{E}}_{2}$ algebra structures.
+However, it was observed by Strickland [STR99, Remark 6.5] at the prime $2$ that, although these forms of truncated Brown–Peterson spectra are equipped with $\operatorname{\mathbb{E}}_{1}$-ring structures, they are not equipped with additional structure.
+In particular, they are not equipped with $\operatorname{\mathbb{E}}_{2}$-ring structures.
 To compute invariants of all $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra we are therefore forced to work without additional structure.
@@
 We therefore resolve the redshift, Lichtenbaum–Quillen and telescope questions in the case of algebraic K-theory of truncated Brown–Peterson spectra.
-Let $T(n)=\mathbb{S}/(p^{i_{0}},\cdots,v_{n-1}^{i_{n-1}})[v_{n}^{-1}]$ for suitably large integers $i_{0},\cdots,i_{n-1}$.
+Let $T(n)=\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}}\cdots,v_{n-1}^{i_{n-1}})[v_{n}^{-1}]$ for suitably large integers $i_{0},i_{1},\cdots,i_{n-1}$.
 We write $L_{n}$ for the Bousfield localization at $\mathbb{Q}\oplus K(1)\oplus\cdots K(n)$ where $K(i)$ is the $i$-th Morava K-theory at the prime $p$.
@@
 Theorem A (Corollary 6.5, Corollary 6.6 and Corollary 6.7).
-Let $p$ be a prime number and $n\geq 1$ an integer.
+Let $p$ be a prime number and $n\geq-1$ be an integer.
 Let $\operatorname{BP}\langle n\rangle$ be an arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of the $n$-th truncated Brown–Peterson spectra at the prime $p$, see Definition 2.1.
@@
   3. (3)
-    (Lichtenbaum–Quillen) The localization map $\operatorname{K}(\operatorname{BP}\langle n\rangle)\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)$ has bounded above fiber.
+    (Lichtenbaum–Quillen) The localization map $\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}$ has bounded above fiber.
 Each of these results follow from the following explicit computation.
 Theorem B (Theorem 5.5 and Corollary 5.7).
-Let $n\geq 1$ be an integer. The mod $(p,v_{1},\cdots,v_{n+1})$ syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
-$\operatorname{\mathbb{F}}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle\oplus\bigoplus_{i=1}^{n+1}\operatorname{\mathbb{F}}_{p}\langle\lambda_{s}:s\in\langle n\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
+Let $n\geq-1$ be an integer. The mod $(p,v_{1},\cdots,v_{n+1})$ syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
+$\operatorname{\mathbb{F}}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle\oplus\bigoplus_{i=1}^{n+1}\operatorname{\mathbb{F}}_{p}\langle\lambda_{s}:s\in\langle n+1\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
 as a bigraded $\operatorname{\mathbb{F}}_{p}$-vector space and the $v_{n+1}$-Bockstein spectral sequence collapses.
-Here the classes are in bidegrees $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ and $\langle n\rangle=\{1,2,\cdots,n\}$.
+Here the classes are in bidegrees $\|\partial\|=(-1,1)$, $\|\lambda_{i}\|=(2p^{i}-1,1)$, $\|\Xi_{j,d}\|=(2p^{j}-1-2dp^{j-1},1)$ and $\langle n+1\rangle=\{1,2,\cdots,n+1\}$.
 Here and throughout we write $\operatorname{\mathbb{F}}_{p}\langle x_{1},x_{2},\cdots,x_{n}\rangle$ for an exterior algebra with generators $x_{1},x_{2},\cdots,x_{n}$.
@@
 Theorem C (Theorem 6.1 and Corollary 6.2).
-Let $p\geq 5$. We can identify
+Let $p\geq 5$ and let $\operatorname{BP}\langle 2\rangle$ be an arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of the second truncated Brown–Peterson spectra at the prime $p$, see Definition 2.1.
+We can identify
 $\operatorname{TC}_{*}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})=\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\partial,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{i=1}^{3}\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\lambda_{s}:s\in\langle 3\rangle-\{i\}\rangle\{\Xi_{i,d}:0<d<p\}$
@@
 Our theorem extends this result to arbitrary $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of truncated Brown–Peterson spectra, such as those arising from the choice of Araki or Hazewinkle generators.
-When $n=0,1$, Theorem B was proven by [LW22] and [HRW22] respectively. As a consequence, Hahn–Raksit–Wilson extend Theorem A and the $n=1$ analogue of Theorem C to the prime $p=3$ in [HRW22].
+When $n=0,1$, Theorem B was proven by [LW22] and [HRW22] respectively. As a consequence, Hahn–Raksit–Wilson [HRW22] extend Theorem A and the $n=1$ analogue of Theorem C to the prime $p=3$.
 When $n=2$, Theorem C was first proven by [AAC+25] at primes $p\geq 7$ for arbitrary $\operatorname{\mathbb{E}}_{3}$ $\operatorname{MU}$-algebra forms.
@@
 In fact, using Theorem B, T. Yang [YAN25, Theorem 2.7] has provided a precise bound in the statement of Theorem A (3), in particular the map
-$\pi_{*}\operatorname{K}(\operatorname{BP}\langle n\rangle)\longrightarrow\pi_{*}L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)$
+$\pi_{*}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}\longrightarrow\pi_{*}L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}$
 is an isomorphism in positive degrees.
@@
 Additionally, the author benefited from input from Christian Ausoni, Robert Burklund, Ishan Levy, Piotr Pstrągowski, Arpon Raksit, John Rognes, Andrew Senger, and Tristan Yang, either from direct communication or indirect communication though Jeremy Hahn and Dylan Wilson.
+The author would also like to thank Shay Ben-Moshe for helpful feedback on a previous draft.
+The author is also grateful to Max Planck Institute for Mathematics in Bonn for its hospitality and financial support.
 2. Hochschild homology
 First, we define forms of $\operatorname{BP}\langle n\rangle$ in Section 2.1 and the $\operatorname{MU}$-based motivic filtration in Section 2.2.
-We compute Hochschild homology of $\operatorname{MU}$, $\operatorname{BP}$, and $\operatorname{BP}\langle n\rangle$ with $\operatorname{\mathbb{F}}_{p}$-coefficients in Section 2.3 along with their motivic filtrations in Section 2.4.
+We compute Hochschild homology of $\operatorname{BP}\langle n\rangle$ with $\operatorname{\mathbb{F}}_{p}$-coefficients in Section 2.3 along with its motivic filtrations in Section 2.4.
 2.1. Forms of $\operatorname{BP}\langle n\rangle$
-By [CM15, Corollary 1.3], the Quillen idempotent defines an $\operatorname{\mathbb{E}}_{2}$ algebra map
+By [CM15, Corollary 1.3], the Quillen idempotent defines an $\operatorname{\mathbb{E}}_{2}$-ring map
 $\operatorname{BP}\to\operatorname{MU}_{(p)}\,.$
-While $\pi_{*}\operatorname{BP}$ is only non-canonically isomorphic to a polynomial ring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots]$, for each $i\geq 1$ the subring $\mathbb{Z}_{(p)}[v_{1},\cdots,v_{i}]\subset\pi_{*}\operatorname{BP}$ is well-defined.
+While $\pi_{*}\operatorname{BP}$ is only non-canonically isomorphic to a polynomial ring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots]$, for each $i\geq 1$ the subring $\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots,v_{i}]\subset\pi_{*}\operatorname{BP}$ is well-defined.
 Indeed, it is the subring generated by all elements of degree at most $2p^{i}-2$.
@@
 An $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle n\rangle$ is a $p$-local $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra $R$ such that the composite
-$\mathbb{Z}_{(p)}[v_{1},\cdots,v_{n}]\subset\operatorname{BP}_{*}\subset\pi_{*}\operatorname{MU}_{(p)}\to\pi_{*}R$
+$\mathbb{Z}_{(p)}[v_{1},v_{2},\cdots,v_{n}]\subset\operatorname{BP}_{*}\subset\pi_{*}\operatorname{MU}_{(p)}\to\pi_{*}R$
 is an isomorphism. The last map in this composite is the $p$-localized unit of the $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra structure.
 In particular, the Eilenberg–MacLane spectrum $\mathbb{Z}_{(p)}$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 0\rangle$ and the Adams summand $\ell$ of $p$-local topological K-theory $\mathrm{ku}$ is a $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle 1\rangle$.
+By convention $\mathbb{F}_{p}$ is a $\mathbb{E}_{\infty}$ $\operatorname{MU}$-algebra form of $\operatorname{BP}\langle-1\rangle$.
 In this paper, we will be interested in the algebraic $K$-theories of $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$.
@@
 We make these choices such that, for each $i>n$, the class $v_{i}$ maps to zero in $\pi_{2p^{i}-2}\operatorname{BP}\langle n\rangle.$
+Convention 2.3.
+Throughout the paper, whenever a prime $p$, the notations $L_{n}$, $L_{n}^{f}$ and $\operatorname{BP}\langle n\rangle$ appear in a statement then the implicit primes are always the same and agree with the given prime $p$ if it appears.
 2.2. The motivic filtration
 We review the definition of the motivic filtration relative to $\operatorname{MU}$, defined on the $\operatorname{THH}$ of any $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra by [HRW22].
-Recollection 2.3.
+Recollection 2.4.
 Suppose that $R$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra. Then we follow [HRW22, Definition 4.2.1] in defining
@@
 In this particular paper, we will always study motivic filtrations relative to $\operatorname{MU}$. Thus, we make the following further simplifying convention:
-Convention 2.4.
+Convention 2.5.
 Given a $\operatorname{THH}(\operatorname{MU})$-module $M$, we abbreviate $\mathrm{fil}^{*}_{\textup{ev}/\operatorname{THH}(\operatorname{MU})}M$ as $\mathrm{fil}^{*}_{\textup{mot}}M$ throughout.
 Therefore, if $R$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra, we write $\mathrm{fil}^{*}_{\textup{mot}}\operatorname{THH}(R)$ for $\mathrm{fil}^{*}_{\textup{mot}/\operatorname{MU}}\operatorname{THH}(R)$.
-Warning 2.5.
+Warning 2.6.
 For a general $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra $R$, using $\mathrm{fil}_{\textup{mot}}^{*}\operatorname{THH}(R)$ to denote $\mathrm{fil}^{*}_{\textup{mot}/\operatorname{MU}}\operatorname{THH}(R)$ may be in contradiction with the notations of [HRW22, PST23].
@@
 $M/v_{i}:=M\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}(\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}/v_{i}).$
-We sometimes write $M/(p,\dots,v_{i})$ for the iterated tensor product
+We sometimes write $M/(p,v_{1},\cdots,v_{i})$ for the iterated tensor product
 $M\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}(\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}/p)\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}\cdots\otimes_{\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}}(\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}/v_{i}).$
@@
 Thus, if $R$ is an $\operatorname{\mathbb{E}}_{1}$ $\operatorname{MU}$-algebra, we may refer to
-$\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(R)/(p,\cdots,v_{i}).$
+$\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(R)/(p,v_{1},\cdots,v_{i}).$
 Remark 2.9.
@@
 sends $v_{i}$ to $v_{i}$. In particular, if $M$ is a $\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}$-module, then
-$M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}}(\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}/(p,\cdots,v_{i}))\cong M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}}(\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}/(p,\cdots,v_{i})).$
+$M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}}(\mathrm{gr}^{*}_{\textup{ev}}\mathbb{S}/(p,v_{1},\cdots,v_{i}))\cong M\otimes_{\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}}(\mathrm{gr}^{*}_{\textup{ev}}\operatorname{MU}/(p,v_{1},\cdots,v_{i})).$
 Definition 2.10.
@@
 Here and throughout given an $\operatorname{\mathbb{E}}_{1}$-ring $R$ and an $R\otimes R^{\textup{op}}$-module $M$ we write $\operatorname{THH}(R;M):=M\otimes_{R\otimes R^{\textup{op}}}R$.
-Recollection 2.10.
+Recollection 2.11.
 There is an isomorphism
@@
 We write $\lambda_{j}\coloneqq\lambda_{p^{j}-1}^{\prime}$.
-The $\operatorname{\mathbb{E}}_{3}$ algebra $\operatorname{THH}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})$ is a retract of $\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$ as an $\mathbb{E}_{1}$ algebra by [CM15, Corollary 1.3] and has homotopy groups
+The $\operatorname{\mathbb{E}}_{3}$-ring $\operatorname{THH}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})$ is a retract of $\operatorname{THH}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$ as an $\mathbb{E}_{1}$-ring by [CM15, Corollary 1.3] and has homotopy groups
 $\operatorname{THH}_{*}(\operatorname{BP};\operatorname{\mathbb{F}}_{p})\cong\mathbb{F}_{p}\langle\lambda_{i}:i\geq 1\rangle\,.$
-Property 2.11.
+Proposition 2.12.
 As an $\operatorname{\mathbb{E}}_{0}$ $\operatorname{THH}(\operatorname{BP})$-algebra,
@@
 where $\|\mu\|=(2,0)$.
-First, we need a preliminary result, which uses the notion of even flatness due to Pstrągowski, see [PST23, Definition 4.4].
-Property 2.12.
-The $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-module $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is an even flat.
+We also say a spectrum $X$ is even if $\pi_{2k-1}X=0$ for all integers $k$. We need a preliminary result, which uses the notion of even flatness due to Pstrągowski, see [PST23, Definition 4.4].
+Proposition 2.13.
+The $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-module $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is even flat.
 In particular, the spectrum $\operatorname{THH}(\operatorname{BP}\langle n\rangle/\operatorname{MU})$ is even.
@@
 $\operatorname{\operatorname{Cotor}}_{(A,\Gamma)}^{*,*}(A,\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}(\mathrm{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})\to\operatorname{\operatorname{Cotor}}_{(A,\Gamma)}^{*,*}(A,\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}(\mathrm{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})$
-of $\operatorname{\mathbb{E}}_{0}$-$\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebras where
+of $\operatorname{\mathbb{E}}_{0}$ $\operatorname{THH}_{*}(\operatorname{MU};\operatorname{\mathbb{F}}_{p})$-algebras where
 $\operatorname{\operatorname{Cotor}}_{(A,\Gamma)}^{*,*}(A,\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{THH}(\mathrm{MU};\operatorname{\mathbb{F}}_{p})}^{\mathbb{L}}\operatorname{\mathbb{F}}_{p})\cong\operatorname{\mathbb{F}}_{p}[\mu]\,.$
@@
 By Proposition 2.13, we know that $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ is even.
-Therefore, the $p,v_{1},\cdots,v_{n}$-Bockstein spectral sequences all collapse and we conclude that $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle/\operatorname{MU})_{p}^{\wedge}$ is a finitely generated free $\operatorname{\mathbb{Z}}_{p}^{\wedge}[v_{1},\cdots v_{n}]$-module.
+Therefore, the $p,v_{1},\cdots,v_{n}$-Bockstein spectral sequences all collapse and we conclude that $\operatorname{THH}_{*}(\operatorname{BP}\langle n\rangle/\operatorname{MU})_{p}^{\wedge}$ is a finitely generated free $\operatorname{\mathbb{Z}}_{p}^{\wedge}[v_{1},v_{2},\cdots,v_{n}]$-module.
 Letting $B_{i}=\operatorname{BP}\langle i\rangle$ for $1\leq i\leq n$.
@@
 for each $0\leq i\leq n$ and after a finite induction we have proven the claim. ∎
+Note that the commutative diagram of $\operatorname{THH}(\operatorname{MU})$-modules
+$\operatorname{MU}$$\operatorname{THH}(\operatorname{MU})$$\operatorname{BP}\langle n\rangle$$\operatorname{THH}(\operatorname{BP}\langle n\rangle)$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{BP}\langle n\rangle;\mathbb{F}_{p})$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$}
+implies that the class $v_{n+1}\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}$ acts trivially on $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\mathbb{F}_{p})$ compatibly with the map $\operatorname{THH}(\operatorname{BP}\langle n\rangle;\mathbb{F}_{p})\to\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$.
+Definition 2.16.
+Consider the short exact sequence
+$0\to M\to M\langle\varepsilon_{n+1}\rangle\to\Sigma^{2p^{n+1}-1,-1}M\to 0$
+where $M=\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})$ induced by the cofiber sequence
+$\Sigma^{2p^{n+1}-2}\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}\overset{v_{n+1}}{\longrightarrow}\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}\to\mathrm{gr}_{\textup{ev}}^{*}\operatorname{MU}/v_{n+1}$
+by base-change and using the observation that $v_{n+1}$ acts trivially on $M$.
+Then the class $\varepsilon_{n+1}$ is defined as a lift of $1\in\Sigma^{2p^{n+1}-1,-1}M$.
+The choice of lift $\varepsilon_{n+1}$ is unique (up to a unit) by Proposition 2.14.
 Theorem 2.17.
@@
 $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n})$
-as well. Taking a further quotient by $v_{n+1}$ produces the result. To see this, consider the commutative diagram of $\operatorname{THH}(\operatorname{MU})$-modules
-$\operatorname{MU}$$\operatorname{THH}(\operatorname{MU})$$\operatorname{BP}\langle n\rangle$$\operatorname{THH}(\operatorname{BP}\langle n\rangle)$$\operatorname{\mathbb{F}}_{p}$$\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$}
-to determine that the $v_{n+1}\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{MU}$ acts trivially on $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)$ compatibly with the map $\operatorname{THH}(\operatorname{BP}\langle n\rangle)\to\operatorname{THH}(\operatorname{\mathbb{F}}_{p})$.
-∎
+as well. Taking a further quotient by $v_{n+1}$ produces the result by Definition 2.16 and the discussion preceding it. ∎
 3. Hodge Tate Cohomology
@@
 Theorem 3.1.
-On bigraded homotopy groups, the commutative diagram
+The commutative diagram
 $\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})\,$$\scriptstyle{\varphi_{p}}$$\scriptstyle{\varphi_{p}}$
@@
 Proof.
-First, note that there is a canonical map
+First, note that by construction there is a map
 $L_{n}^{f}\operatorname{MU}\longrightarrow L_{n}^{f}\operatorname{BP}\langle n\rangle$
@@
 $\sigma^{2}v_{n+1}=\mu^{p^{n+1}}\in\pi_{2p^{n+1}}\operatorname{THH}(\operatorname{\mathbb{F}}_{p}/\operatorname{MU})\,.$
-This produces a element $\mu^{p^{n+1}}$ in the bigraded commutative commutative ring
+This produces an element $\mu^{p^{n+1}}$ in the bigraded commutative commutative ring
 $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\tau_{\geq 0}L_{n}^{f}\operatorname{MU})$
@@
 $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p})\longrightarrow\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle;\operatorname{\mathbb{F}}_{p}^{\otimes p})^{tC_{p}}$
-as a map of $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]$-algebras and it makes sense to ask whether the map is given by inverting the class $\mu^{p^{n}}$.
+as a map of $\operatorname{\mathbb{F}}_{p}[\mu^{p^{n}}]$-modules and it makes sense to ask whether the map is given by inverting the class $\mu^{p^{n}}$.
 As in [HW22, Recollection 4.3.2], we can determine that the associated graded $\mathrm{gr}_{\textup{Ad}}^{*}\operatorname{BP}\langle n\rangle$ of the filtration
@@
 $\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})\otimes_{\operatorname{\mathbb{F}}_{p}}\pi_{*}\operatorname{\mathbb{F}}_{p}\otimes\operatorname{THH}(\operatorname{\mathbb{S}}[v_{0},\dots,v_{n}];\operatorname{\mathbb{S}})$$\operatorname{THH}_{*}(\operatorname{\mathbb{F}}_{p})^{tC_{p}}\otimes_{\operatorname{\mathbb{F}}_{p}}\pi_{*}\operatorname{\mathbb{F}}_{p}\otimes\operatorname{THH}(\operatorname{\mathbb{S}}[v_{0},\dots,v_{n}];\operatorname{\mathbb{S}}^{\otimes p})^{tC_{p}}$}
-and in fact the same considerations as in the proof of [HW22, Theorem 4.3.1] allow us to determine that the map is given by inverting $\mu$ on the $\operatorname{\mathbb{E}}_{2}$-pages.
-We have completely computed the top spectral sequence and since the bottom spectral sequence is given by inverting $\mu$ on $\operatorname{\mathbb{E}}_{2}$-pages, we determine that at each page the spectral sequence is given by inverting some power of $\mu$, and hence remains a monomomorphism entirely determining also the spectral sequence on the bottom.
+and in fact the same considerations as in the proof of [HW22, Theorem 4.3.1] allow us to determine that the map is given by inverting $\mu$ on the $\mathrm{E}_{2}$-pages.
+We have completely computed the top spectral sequence and since the bottom spectral sequence is given by inverting $\mu$ on $\mathrm{E}_{2}$-pages, we determine that at each page the spectral sequence is given by inverting some power of $\mu$, and hence remains a monomomorphism entirely determining also the spectral sequence on the bottom.
 This proves the claim.
 ∎
-Theorem 3.2.
+Proposition 3.2.
 The spectral sequence beginning with
 $\left(\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})\right)[t]$
-and converging to $\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$ collapses after the $d_{1}$ differential.
+and converging to $\pi_{*}\mathrm{gr}^{*}_{\textup{mot}}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$ (up to isomorphism) collapses after the $d_{1}$ differential.
 All permanent cycles are on the 0-line.
 Proof.
-Since $v_{n+1}$ is detected by $t\mu^{p^{n+1}}$, we determine that $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$.
-Using the action of $\lambda_{1},\cdots,\lambda_{n+1}$ together with the monomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\cdots,\lambda_{n+1})$, we determine a Leibniz rule $d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$ for all
+Since $v_{n+1}$ is detected by $t\mu^{p^{n+1}}$, we determine that $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ by Definition 2.16 ($\varepsilon_{n+1}$ corresponds to a null homotopy of $v_{n+1}$).
+Using the action of $\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}$ together with the monomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1})$, we determine a Leibniz rule $d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$ for all
 $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n})\,.$
-Since $\mu$ is a unit in a ring that acts on this spectral sequence, mainly the corresponding spectral sequence for $\tau_{\geq 0}L_{n}^{f}\operatorname{MU}$, the spectral sequence collapses after running the first differential and all permanent cycles are concentrated on the 0-line.
+Since $\mu$ is a unit in an algebra spectral sequence that acts on this spectral sequence, namely the corresponding spectral sequence for $\tau_{\geq 0}L_{n}^{f}\operatorname{MU}$ (cf.
+proof of Theorem 3.1), the spectral sequence collapses after running the first differential and all permanent cycles are concentrated on the 0-line.
 Moreover, we can conclude that
@@
 $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{\mathbb{F}}_{p})/(p,v_{1},\cdots,v_{n+1})$
-of $\operatorname{\mathbb{E}}_{0}$-$\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})$-algebras induced by the canonical $\operatorname{\mathbb{E}}_{1}$-$\operatorname{BP}$-algebra reduction map
+of $\operatorname{\mathbb{E}}_{0}$ $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP})$-algebras induced by the canonical $\operatorname{\mathbb{E}}_{1}$ $\operatorname{BP}$-algebra reduction map
 $\operatorname{BP}\langle n\rangle\to\tau_{\leq 0}\operatorname{BP}\langle n\rangle=\operatorname{\mathbb{Z}}_{(p)}\to\operatorname{\mathbb{F}}_{p}.$
@@
 $d_{p^{m}}(t^{jp^{m-1}}\lambda_{S})\dot{=}t^{p^{m}+jp^{m-1}}\lambda_{m}\lambda_{S}$
-for $\ell\in\mathbb{Z}$, $0<j<p$ and $S,T\subset\{1,\cdots,n+1\}$ with $m\not\in S$. The spectral sequence then collapses at the $p^{n+1}+1$-page without room for further differentials. Consequently,
+for $0<j<p$ and $S\subset\{1,\cdots,n+1\}-\{m\}$. The spectral sequence then collapses at the $p^{n+1}+1$-page without room for further differentials. Consequently,
 $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\cong\mathbb{F}_{p}\langle\lambda_{1},\dots,\lambda_{n+1}\rangle[t^{\pm p^{n+1}}]$
@@
 Proof.
-The differential $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ follows from the definition of $\varepsilon_{n+1}$ and Theorem 2.15. The Leibniz rule
+The differential $d_{1}(\varepsilon_{n+1})=t\mu^{p^{n+1}}$ follows from Definition 2.16 ($\varepsilon_{n+1}$ corresponds to a null homotopy of $v_{n+1}$) and Theorem 2.17. The Leibniz rule
 $d_{1}(x\varepsilon_{n+1})=xt\mu^{p^{n+1}}$
-for all $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})[t^{\pm 1}]$ follows from [AHW24b, Corollary 3.2.6], which implies that the classes $\lambda_{1},\cdots,\lambda_{n+1}$ are permanent cycles that act on this spectral sequence, together with the monomomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\cdots,\lambda_{n+1})$.
+for all $x\in\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n})[t^{\pm 1}]$ follows from [AHW24b, Corollary 3.2.6], which implies that the classes $\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}$ are permanent cycles that act on this spectral sequence, together with the monomomorphism into the spectral sequence for $\operatorname{\mathbb{F}}_{p}$ modulo $(\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1})$.
 The differentials
@@
 for all $S\subset\{1,\cdots,n+1\}$ with $m\not\in S$ for each $m\geq 1$ follow from the action of the spectral sequence for $\operatorname{BP}$ and the differentials in that spectral sequence, computed in [AHW24b, Proposition 3.2.9].
-Finally, we simply check bidegrees and observe that the spectral sequence must collapse at the $\mathrm{E}_{p^{n+1}+1}$-page.
+By [AHW24b, Proposition 3.2.9], we also know that that $t^{p^{n+1}}$ survives to the $\mathrm{E}_{p^{n+2}}$-page and it acts on the spectral sequence up until this page, so $\mathrm{E}_{p^{n+1}+1}=\mathrm{E}_{p^{n+2}}$.
+Finally, we simply check bidegrees and observe that the spectral sequence must collapse at the $\mathrm{E}_{p^{n+2}}$-page.
 ∎
+Remark 4.1.
+Note that, as a consequence of Proposition 4.2, we have an isomorphism
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\cong\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n})\,.$
+This could alternatively be proven directly along the lines of [HRW22, § 6.4] and we believe that such a result holds quite generally, however one still needs to understand the Nygaard filtration on prismatic cohomology to compute syntomic cohomology and this is also achieved in Proposition 4.2
 5. Syntomic cohomology
@@
 $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\,.$
-We note that these bigraded groups are equipped with the Nygaard filtration coming the $t$-Bockstein spectral sequence and therefore there is an exact sequence
+We note that these bigraded groups are equipped with the Nygaard filtration coming from the $t$-Bockstein spectral sequence and therefore there is an exact sequence
 $0\to\mathrm{Nyg}_{\geq 1}\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\mathrm{Nyg}_{=0}\to 0$
@@
 where
-$A_{10}$ $:=\mathrm{Nyg}_{\geq 1}\cap T\,$,
-$A_{00}$ $:=\mathrm{Nyg}_{\geq 1}/A_{10}\,$,
-$A_{11}$ $:=T/A_{10}\text{ and }$
-$A_{10}$ $=F/A_{00}\,.$
-Property 5.2.
+$A_{11}$ $:=\mathrm{Nyg}_{\geq 1}\cap T\,$,
+$A_{10}$ $:=\mathrm{Nyg}_{\geq 1}/A_{11}\,$,
+$A_{01}$ $:=T/A_{11}\text{ and }$
+$A_{00}$ $=F/A_{10}\,.$
+Proposition 5.2.
 The classes in the $\mathrm{E}_{2}$-page of the periodic $t$-Bockstein spectral sequence that are hit by differentials that cross from negative Nygaard filtration to positive Nygaard filtration are exactly
-$A_{10}:=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,.$
+$A_{11}=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,.$
 Proof.
 This follows from a careful bookkeeping of the differentials determined in Proposition 4.2. ∎
-The following corollary is immediate.
+The following corollary is immediate from Proposition 4.2 and Proposition 5.2.
 Corollary 5.3.
 We compute that
-$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{p^{n+1}}]\oplus$
-$\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{d,j}:0<d<p\}$
-as a $\mathbb{F}_{p}\langle\lambda_{1},\cdots\lambda_{n+1}\rangle$-module where $\Xi_{d,j}$ is detected by $t^{dp^{j-1}}\lambda_{j}$, $y^{p^{n+1}}$ is detected by $t^{-p^{n+1}}$ and we abuse notation and write $\lambda_{k}$ for the class detected by $\lambda_{k}$.
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=A_{00}\oplus A_{10}\oplus A_{01}\oplus A_{11}$
+as $\mathbb{F}_{p}\langle\lambda_{1},\lambda_{2},\cdots\lambda_{n+1}\rangle$-modules where
+$A_{00}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
+$A_{11}$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{\Xi_{d,j}:0<d<p\}\,$,
+$A_{01}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]\{\mu^{p^{n+1}}\}\text{ and }$
+$A_{10}$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[t^{p^{n+1}}]\{t^{p^{n+1}}\}\,.$
+Here $\Xi_{d,j}$ is detected by $t^{dp^{j-1}}\lambda_{j}$, we abuse notation and write $t^{p^{n+1}}$ for the class detected by $t^{p^{n+1}}$, we write $\lambda_{k}$ for the class detected by $\lambda_{k}$ and we write $\mu^{kp^{n+1}}$ for the class detected by $\mu^{kp^{n+1}}$.
 Moreover, the canonical map
 $\textup{can}:\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
-sends $\lambda_{k}$ to $\lambda_{k}$, $y^{p^{n+1}}$ to $y^{p^{n+1}}$, and sends $\lambda_{s}^{\epsilon}\cdot\Xi_{d,j}$ to zero for each $\epsilon\in\{0,1\}$, $s\in\{1,\cdots,n\}-\{j\}$, and $j\in\{1,\cdots,n+1\}$.
+sends $\lambda_{k}$ to $\lambda_{k}$, $t^{p^{n+1}}$ to $t^{p^{n+1}}$, and sends $\lambda_{S}\cdot\Xi_{d,j}$ and $\lambda_{T}\mu^{kp^{n+1}}$ to zero for each $k\geq 1$, $S\subset\{1,\cdots,n+1\}-\{j\}$, $j\in\{1,\cdots,n+1\}$, $0<d<p$ and $T\subset\{1,\cdots,n+1\}$.
+In other words, $\mathrm{can}|_{A_{01}\oplus A_{11}}=0$ and $\mathrm{can}|_{A_{00}\oplus A_{10}}=\mathrm{inc}_{A_{00}\oplus A_{10}}$ where $\textup{inc}_{A_{00}\oplus A_{10}}$ is the canonical inclusion.
 We further need to compute the Frobenius map. We proceed by considering the map of $t$-Bockstein spectral sequences
 $\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})[t]$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,v_{1},\cdots,v_{n+1})[t]$$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\,.$}
-Property 5.4.
+Proposition 5.4.
 The $t$-Bockstein spectral sequence
-$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n+1})[t]\implies\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}/(p,\cdots,v_{n+1})[t]\implies\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}})^{hS^{1}}/(p,v_{1},\cdots,v_{n+1})$
 has $\mathrm{E}_{2}$-term
-$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\otimes\mathbb{F}[\mu^{\pm p^{n}},t]\langle\varepsilon_{n+1}\rangle$
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}}=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{\pm p^{n}},t]\langle\varepsilon_{n+1}\rangle$
 and it collapses after the differential
@@
 where $x\in\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{\pm p^{n}},t]$. We can therefore identify
-$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{\pm p^{n+1}}]$
-where $y^{p^{n+1}}$ is detected by $\mu^{p^{n+1}}$.
-Letting
-$A$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
-$B$ $=\bigoplus_{j=1}^{n+1}\bigotimes_{s\in\{1,\cdots,n+1\}-\{j\}}\mathbb{F}_{p}\langle\lambda_{s}\rangle\{\Xi_{j,d}:0<d<p\}\,,\text{ and }$
-$C$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{p^{n+1}}]\{y^{p^{n+1}}\}$
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}(\operatorname{THH}(\operatorname{BP}\langle n\rangle)^{tC_{p}})^{hS^{1}}/(p,v_{1},\cdots,v_{n+1})=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{\pm p^{n+1}}]$
+where $y^{\pm kp^{n+1}}$ is detected by $\mu^{\pm kp^{n+1}}$.
 The Frobenius map
 $\varphi:\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})\to\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
-is given by
-$\varphi|_{A}=\textup{inc}_{A}\circ\mathrm{id}_{A}$
-and $\varphi|_{B\oplus C}=0$ where $\textup{inc}_{A}:A\to A\oplus B\oplus C$ is the canonical inclusion.
+is given by $\varphi|_{A_{00}}=\textup{inc}_{A_{00}}$, $\varphi|_{A_{01}}$ is an isomorphism and $\varphi|_{A_{10}\oplus A_{11}}=0$, where $\textup{inc}_{A_{00}}$ is the canonical inclusion.
+Proof.
+This follows from Proposition 3.2 and Corollary 5.3 together with [AHW24b, Corollary 3.2.6], which implies that the classes in $\mathbb{F}_{p}\langle\lambda_{1},\lambda_{2},\cdots,\lambda_{n+1}\rangle$ are equalized by $\mathrm{can}$ and $\varphi$.
+∎
 We now present our computation of syntomic cohomology of $\operatorname{BP}\langle n\rangle$.
 Theorem 5.5.
-The mod $(p,\dots,v_{n+1})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
-(5.6) $\mathbb{F}_{p}\langle\delta,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
+Let $n\geq-1$. The mod $(p,v_{1},\dots,v_{n+1})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
+(5.6) $\mathbb{F}_{p}\langle\partial,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
 as a $\mathbb{F}_{p}\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module.
@@
 Let
-$A$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
-$B$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}\,$,
-$C$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[y^{-p^{n+1}}]\{y^{-p^{n+1}}\}$
-then we determined that $\mathrm{can}|_{A}=\textup{incl}_{A\oplus C}\circ\mathrm{id}_{A\oplus C}$ and $\mathrm{can}(B)=0$ in Corollary 5.3, where $\textup{incl}_{A\oplus C}$ denotes the canonical inclusion $A\oplus C\to A\oplus B\oplus C$.
-We determined that $\varphi|_{A}=\textup{inc}_{A}\circ\mathrm{id}_{A}$ and $\varphi|_{B\oplus C}=0$ in Proposition 5.4.
-We therefore determine that $\ker(\mathrm{can}-\varphi)=A\oplus B$ and $\operatorname{coker}(\mathrm{can}-\varphi)=A$.
-We write $\partial$ to indicate that these classes come from the boundary map in the long exact sequence mod $(p,v_{1},\cdots,v_{n+1})$ induced by the fiber sequence
+$A_{00}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle\,$,
+$A_{10}$ $=\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\otimes\operatorname{\mathbb{F}}_{p}\{t^{dp^{j-1}}\lambda_{j}:0<d<p\}\,$,
+$A_{01}$ $=\operatorname{\mathbb{F}}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[\mu^{p^{n+1}}]\{\mu^{p^{n+1}}\}\text{ and }$
+$A_{11}$ $=\mathbb{F}_{p}\langle\lambda_{1},\cdots,\lambda_{n+1}\rangle[t^{p^{n+1}}]\{t^{p^{n+1}}\}\,.$
+We determined that $\mathrm{can}|_{A_{01}\oplus A_{11}}=0$ and $\mathrm{can}|_{A_{00}\oplus A_{10}}=\mathrm{inc}_{A_{00}\oplus A_{10}}$ in Corollary 5.3.
+We determined that $\varphi|_{A_{00}}=\textup{inc}_{A_{00}}$, $\varphi|_{A_{01}}$ is an isomorphism and $\varphi|_{A_{10}\oplus A_{11}}=0$ in Proposition 5.4.
+We therefore conclude that $\ker(\mathrm{can}-\varphi)=A_{00}\oplus A_{11}$ and $\operatorname{coker}(\mathrm{can}-\varphi)=A_{00}$.
+We write $\partial$ to indicate that classes that are multiples $\partial$ of come from the boundary map in the long exact sequence mod $(p,v_{1},\cdots,v_{n+1})$ induced by the fiber sequence
 $\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)\to\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}^{-}(\operatorname{BP}\langle n\rangle)\overset{\mathrm{can}-\varphi}{\longrightarrow}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TP}(\operatorname{BP}\langle n\rangle)\,.$
@@
 Corollary 5.7.
-The mod $(p,\dots,v_{n})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
-(5.8) $\mathbb{F}_{p}[v_{n+1}]\langle\delta,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}[v_{n+1}]\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}.$
+Let $n\geq-1$. The mod $(p,\dots,v_{n})$-syntomic cohomology of $\operatorname{BP}\langle n\rangle$ is
+(5.8) $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\dots\lambda_{n+1}\rangle\oplus\bigoplus_{j=1}^{n+1}\mathbb{F}_{p}[v_{n+1}]\langle\lambda_{s}:1\leq s\leq n+1,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}.$
 In particular, it is a bigraded free $\operatorname{\mathbb{F}}_{p}[v_{n+1}]$-module of dimension $2^{n+2}+2^{n}(n+1)(p-1)$ with generators concentrated in degrees $1,\sum_{i=1}^{n+1}2p^{i}-n-1]$ and Adams weights $[0,n+2]$.
@@
 for all primes $p>2$.
-This handles all the $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module generators and therefore proves the claim and therefore the $v_{n+1}$, $\partial$, $\lambda_{1}$, $\cdots$ $\lambda_{n+1}$-linearity proves the claim.
+This handles all the $\mathbb{F}_{p}[v_{n+1}]\langle\partial,\lambda_{1},\cdots,\lambda_{n+1}\rangle$-module generators and therefore the $v_{n+1}$, $\partial$, $\lambda_{1}$, $\cdots$ $\lambda_{n+1}$-linearity proves the claim.
 ∎
-Remark 5.9.
-The dimension of the syntomic cohomology of $\operatorname{BP}\langle n\rangle$ mod $(p,v_{1},\cdots,v_{n+1})$ was predicted by Ausoni–Rognes [AR02, p.
-7], as stated precisely in [ROG01, §2].
-They also claim that when $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{\infty}$-ring and a Smith–Toda complex $\operatorname{\mathbb{S}}/(p,v_{1},\cdots,v_{n})$ exists, then the $v_{n+1}$-Bockstein spectral sequence and the motivic spectral sequence collapse.
-It follows from [AAC+25], that at least when $p\geq 7$ and $n=2$ it suffices to know that $\operatorname{BP}\langle n\rangle$ is an $\operatorname{\mathbb{E}}_{3}$-ring, which is known by [HW22].
 6. Algebraic K-theory
-In this section, we compute algebraic K-theory of $\operatorname{BP}\langle n\rangle$ at sufficiently low heights.
-We then prove Lichtenbaum—Quillen properties for algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
-Finally, we prove that the telescope conjecture is true for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
+In this section, we compute algebraic K-theory of $\operatorname{BP}\langle 2\rangle$ at primes $p\geq 5$.
+We then prove the Lichtenbaum—Quillen property for algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
+Finally, we prove that the telescope conjecture holds for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$.
 6.1. Algebraic K-theory at low heights
@@
 Let $p\geq 5$. Then $\pi_{*}\operatorname{TC}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})$ is isomorphic to
-$\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}[v_{3}]\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
-as $\mathbb{F}_{p}[v_{3}^{d}]$-modules for some integer $d\geq 1$.
+$\operatorname{\mathbb{F}}_{p}[v_{3}]\langle\partial,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}[v_{3}]\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}$
+as $\mathbb{F}_{p}[v_{3}^{d}]\langle\partial,\lambda_{1},\lambda_{2},\lambda_{3}\rangle$-modules for some integer $d\geq 1$.
 Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
@@
 Corollary 6.2.
-Let $p\geq 5$. We compute
+Let $p\geq 5$. There is a preferred isomorphism
 $\pi_{s}\mathrm{K}(\operatorname{BP}\langle 2\rangle)/(p,v_{1},v_{2})\cong\begin{cases}0&s<0\\ \pi_{s}\mathrm{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})\oplus\mathbb{F}_{p}\{\Sigma^{-2}\overline{\tau}_{i}\}&s=2p^{i}-3,i=1,2\\ \pi_{s}\mathrm{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})\oplus\mathbb{F}_{p}\{\Sigma^{-2}\overline{\tau}_{1}\overline{\tau}_{2}\}&s=2p^{2}+2p-4\\ \pi_{s}\mathrm{TC}(\operatorname{BP}\langle 2\rangle/(p,v_{1},v_{2})&\text{ otherwise }\end{cases}$
@@
 $\operatorname{\mathbb{F}}_{p}[v_{3}^{\pm 1}]\otimes\bigl(\mathbb{F}_{p}\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle\oplus\bigoplus_{j=1}^{3}\mathbb{F}_{p}\langle\lambda_{s}:1\leq s\leq 3,s\neq j\rangle\{\Xi_{j,d}:0<d<p\}\bigr)$
-as $\mathbb{F}_{p}[v_{3}^{\pm d}]$-modules for some integer $d\geq 1$.
+as $\mathbb{F}_{p}[v_{3}^{\pm d}]\langle\delta,\lambda_{1},\lambda_{2},\lambda_{3}\rangle$-modules for some integer $d\geq 1$.
 Here $|v_{3}|=2p^{3}-2$, $|\partial|=-1$, $|\lambda_{i}|=2p^{i}-1$ and $|\Xi_{j,d}|=2p^{j}-1-2dp^{j-1}$ for all $1\leq i,j\leq n+1$ and $0<d<p$.
@@
 Theorem 6.4.
-For any integers $i_{0},\cdots,i_{n+1}$ such that the generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$ exists, the graded abelian group
-$\pi_{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
+For any integers $i_{0},i_{1},\cdots,i_{n+1}$ such that the generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ exists, the graded abelian group
+$\pi_{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
 is finite; i.e.
@@
 By Theorem 5.5, the bigraded abelian group
-$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p,\cdots,v_{n+1})$
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p,v_{1},\cdots,v_{n+1})$
 is finite in the sense it is finite in all bidegrees and only non-trivial in finitely many bidegrees. This immediately implies that
-$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
-is finite for any integers $i_{0},\cdots,i_{n+1}$.
-Choosing integers $i_{0},\cdots,i_{n+1}$ such that a generalized Smith–Toda complex $\mathbb{S}/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$ exists, this immediately implies that
-$\pi_{*}\operatorname{TC}(R)/(p^{i_{0}},\cdots,v_{n+1}^{i_{n+1}})$
+$\pi_{*}\mathrm{gr}_{\textup{mot}}^{*}\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
+is finite for any integers $i_{0},i_{1},\cdots,i_{n+1}$ using the finite $p,v_{1},\cdots,v_{n+1}$-Bockstein filtrations.
+Choosing integers $i_{0},i_{1},\cdots,i_{n+1}$ such that the Moore spectrum $\mathbb{S}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ exists, this immediately implies that
+$\pi_{*}\operatorname{TC}(R)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
 is finite by motivic spectral sequence and consequently $\operatorname{TC}(R)$ has fp-type $n+1$ in the sense of [MR99]. By [DGM13, Theorem 3.1.14], there is a fiber sequence
 $\operatorname{K}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})\to\operatorname{TC}(\operatorname{BP}\langle n\rangle)/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})\to\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$
-the algebraic K-theory of $R$ has fp-type $n+1$ since $\pi_{*}\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ is also finite. ∎
+and the algebraic K-theory of $R$ has fp-type $n+1$ since $\pi_{*}\mathbb{Z}/(p^{i_{0}},v_{1}^{i_{1}},\cdots,v_{n+1}^{i_{n+1}})$ is also finite. ∎
 Remark 6.1.
 This extends [HW22, Theorem B] to arbitrary $\mathbb{E}_{1}$ $\operatorname{MU}$-algebra forms of $\operatorname{BP}\langle n\rangle$.
-This implies the higher height analogue of the Lichtenbaum–Qullen conjecture as proposed by Ausoni and Rognes in [AR08].
+This implies the higher height analogue of the Lichtenbaum–Qullen conjecture as proposed by Ausoni and Rognes in [AR08] as part of the family of conjectures known as the redshift conjectures.
 Corollary 6.5.
-The the fiber of the map
+Let $n\geq-1$ and $p$ be a prime. The fiber of the map
 $\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}\longrightarrow L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)_{(p)}$
@@
 Proof.
-This follows from Theorem 6.4 by [MR99, Theorem 8.2(2)] as observed in [HW22, Theorem 3.1.3]. ∎
+This follows from Theorem 6.4 by [MR99, Theorem 8.2(2)] as observed in [HW22, Theorem 3.1.3] together with an arithmetic fracture square argument. ∎
 6.3. Telescope
-We now note that our results also imply the telescope conjecture for topological cyclic homology of $\operatorname{BP}\langle n\rangle$.
+Since it is now known that the telescope conjecture is false in general [BHL+23], it becomes an interesting question to ask for which spectra $X$ is the localization map
+$L_{n+1}^{f}X\to L_{n+1}X$
+is an equivalence.
+For example, it was conjectured by [MR99, Conjecture 3.9] that the telescope conjecture holds for $fp$-spectra.
+We show that our results imply the telescope conjecture for the algebraic K-theory of $\operatorname{BP}\langle n\rangle$, giving some support for this conjecture in light of Theorem 6.4, which implies that the algebraic K-theory of $\operatorname{BP}\langle n\rangle$ is an $fp$-spectrum.
 Corollary 6.6.
-The localization map
+Let $n\geq-1$. The localization map
 $L_{n+1}^{f}\operatorname{K}(\operatorname{BP}\langle n\rangle)\to L_{n+1}\operatorname{K}(\operatorname{BP}\langle n\rangle)$